### v3

# Smooth Tchebycheff Scalarization for Multi-Objective Optimization

Abstract
Multi-objective optimization problems can be found in many real-world applications, where the objectives often conflict each other and cannot be optimized by a single solution. In the past few decades, numerous methods have been proposed to find Pareto solutions that represent optimal trade-offs among the objectives for a given problem. However, these existing methods could have high computational complexity or may not have good theoretical properties for solving a general differentiable multi-objective optimization problem. In this work, by leveraging the smooth optimization technique, we propose a lightweight and efficient smooth Tchebycheff scalarization approach for gradient-based multi-objective optimization. It has good theoretical properties for finding all Pareto solutions with valid trade-off preferences, while enjoying significantly lower computational complexity compared to other methods. Experimental results on various real-world application problems fully demonstrate the effectiveness of our proposed method.
[…]
1 Introduction
Many real-world applications involve multi-objective optimization problems, such as learning an accurate but also fair model (Martinez et al., 2020), building a powerful yet energy-efficient agent (Xu et al., 2020), and finding a drug design that satisfies various criteria (Xie et al., 2021). However, these objectives often conflict each other, making it impossible to optimize them simultaneously with a single solution. For each nontrivial multi-objective optimization problem, there exists a Pareto set that probably contains an infinite number of Pareto solutions that represent different optimal trade-offs among the objectives (Miettinen, 1999; Ehrgott, 2005). Numerous algorithms have been proposed to find a single solution or a finite set of solutions to approximate the Pareto set. This work focuses on the gradient-based methods for differentiable multi-objective optimization.
There are two main types of gradient-based multi-objective optimization algorithms: the scalarization approach (Miettinen, 1999; Ehrgott, 2005) and the adaptive gradient method (Fliege et al., 2019). However, both algorithms have their own disadvantages. The simple linear scalarization is the most straightforward method, but will completely miss all the solutions on the non-convex part of the optimal Pareto front (Das & Dennis, 1997; Ehrgott, 2005). The classical Tchebycheff scalarization can find all Pareto solutions that satisfy the decision-maker’s exact trade-off preference (Bowman, 1976; Steuer & Choo, 1983), but it suffers from a slow convergence rate due to its nonsmooth (i.e., non-differentiable) nature. To overcome the limitations of scalarization approaches, many adaptive gradient methods have been proposed in the past few decades (Fliege & Svaiter, 2000; Schäffler et al., 2002; Désidéri, 2012). These methods aim to find a valid gradient direction that improves the performance of all objectives at each iteration. While these methods have theoretical guarantees for finding a Pareto stationary solution, their high pre-iteration computational complexity makes them less suitable for handling large-scale problems in real-world applications.
Table 1: Comparison of different methods for gradient-based multi-objective optimization.
Required # Iterations Pre-Iteration Complexity Non-convex Pareto Front
Adaptive Gradient Method Small High Yes
Linear Scalarization Small Low No
Tchebycheff (TCH) Scalarization Large Low Yes
STCH Scalarization (This Work) Small Low Yes
[…]
Figure 1: Different multi-objective optimization methods. (a) The Pareto Front is the achievable boundary of the feasible region that represents different (maybe infinite) optimal trade-offs among the objectives. (b) Adaptive Gradient Algorithm aims to find a valid gradient direction to improve the performance of all objectives, which involves solving a quadratic programming problem at each iteration. (c) Linear scalarization cannot find any Pareto solution on the non-convex part of the Pareto front, of which those solutions do not have supporting hyperplanes. (d) Tchebycheff (TCH) Scalarization is capable of finding all Pareto solutions, but requires a large number of iterations. (e) Smooth Tchebycheff (STCH) Scalarization proposed in this work can find all Pareto solutions under mild conditions, while enjoying a much faster convergence speed.
Instead of proposing another adaptive gradient method, this work revisits the straightforward scalarization approach. In particular, we leverage the powerful smooth approximation technique (Nesterov, 2005; Beck & Teboulle, 2012) to develop a lightweight yet efficient scalarization approach with promising theoretical properties. The main contributions of this work are summarized as follows.  [^1] Our source code is available at: github.com/Xi-L/STCH.
  • We propose a smooth Tchebycheff (STCH) scalarization approach for gradient-based multi-objective optimization, which can serve as a fast alternative to the classic Tchebycheff scalarization.
  • We provide detailed theoretical analyses to demonstrate that STCH scalarization has promising theoretical properties while enjoying low computational complexity.
  • We further generalize STCH scalarization to support efficient Pareto set learning.
  • We conduct various experiments on diverse multi-objective optimization problems. The results confirm the effectiveness of our proposed STCH scalarization.
[…]
Figure 2: The advantage of our proposed smooth Tchebycheff scalarization. (a) Problem & Target: We want to find a Pareto solution with an exact trade-off $\bm{\lambda}=(0.5,0.5)$ on the Pareto front. (b) Classical Tchebycheff (TCH) scalarization suffers from a slow convergence speed with an oscillation trajectory. (c) Our proposed smooth Tchebycheff (STCH) scalarization quickly converges to the exact target solution with a smooth trajectory. (d) & (e) The mean/median gaps v.s. number of function evaluations of different methods to the target objective value with $100$ trials.
2 Preliminaries and Related Work for Multi-Objective Optimization
In this work, we consider the following continuous multi-objective optimization problem (MOP) with $m$ differentiable objective functions $\bm{f}=\{f_{1},\cdots,f_{m}\}:\mathcal{X}\rightarrow\mathbb{R}^{m}$ in the decision space $\mathcal{X}\subseteq\mathbb{R}^{n}$:
$\min_{\bm{x}\in\mathcal{X}}\ \bm{f}(\bm{x})=(f_{1}(\bm{x}),f_{2}(\bm{x}),\cdots,f_{m}(\bm{x})).$ (1)
For a non-trivial problem (1), the objective functions $\{f_{1},\cdots,f_{m}\}$ will conflict each other and cannot be simultaneously optimized by a single best solution. For multi-objective optimization, we have the following definition on dominance relation among solutions (Miettinen, 1999):
Definition 2.1 (Dominance and Strict Dominance).
Let $\bm{x}^{(a)},\bm{x}^{(b)}\in\mathcal{X}$ be two solutions for problem (1), $\bm{x}^{(a)}$ is said to dominate $\bm{x}^{(b)}$, denoted as $\bm{f}(\bm{x}^{(a)})\prec\bm{f}(\bm{x}^{(b)})$, if and only if $f_{i}(\bm{x}^{(a)})\leq f_{i}(\bm{x}^{(b)})\ \forall i\in\{1,...,m\}$ and $f_{j}(\bm{x}^{(a)})<f_{j}(\bm{x}^{(b)})\ \exists j\in\{1,...,m\}$. In addition, $\bm{x}^{(a)}$ is said to strictly dominate $\bm{x}^{(b)}$ (i.e., $\bm{f}(\bm{x}^{(a)})\prec_{\text{strict}}\bm{f}(\bm{x}^{(b)})$), if and only if $f_{i}(\bm{x}^{(a)})<f_{i}(\bm{x}^{(b)})\ \forall i\in\{1,...,m\}$.
The (strict) dominance relation only establishes a partial order among different solutions since any two solutions may be non-dominated with each other (e.g., $\bm{f}(\bm{x}^{(a)})\nprec\bm{f}(\bm{x}^{(b)})$ and $\bm{f}(\bm{x}^{(b)})\nprec\bm{f}(\bm{x}^{(a)})$). In this case, the solution $\bm{x}^{(a)}$ and $\bm{x}^{(b)}$ are said to be incomparable. Therefore, we have the following definition of Pareto optimality to describe the optimal solutions for multi-objective optimization:
Definition 2.2 ((Weakly) Pareto Optimality).
A solution $\bm{x}^{\ast}\in\mathcal{X}$ is Pareto optimal if there is no $\bm{x}\in\mathcal{X}$ such that $\bm{f}(\bm{x})\prec\bm{f}(\bm{x}^{\ast})$. A solution $\bm{x}^{\prime}\in\mathcal{X}$ is weakly Pareto optimal if there is no $\bm{x}\in\mathcal{X}$ such that $\bm{f}(\bm{x})\prec_{\text{strict}}\bm{f}(\bm{x}^{\prime})$.
Similarly, a solution $\bm{x}^{*}$ is called local (weakly) Pareto optimal if it is (weakly) Pareto optimal in $\mathcal{X}\cap B(\bm{x}^{*},\delta)$, of which $B(\bm{x}^{*},\delta)=\left\{\bm{x}\in\mathbb{R}^{n}|||x-x^{*}||<\delta\right\}$ is an open ball around $\bm{x}^{*}$ with radius $\delta>0$. In general, the Pareto optimal solution is not unique, and the set of all Pareto optimal solutions is called the Pareto set:
$\bm{X}^{*}=\{\bm{x}\in\mathcal{X}|\bm{f}(\hat{}\bm{x})\nprec\bm{f}(\bm{x})\ \forall\hat{}\bm{x}\in\mathcal{X}\}.$ (2)
Its image in the objective space is called the Pareto front:
$\bm{f}(\bm{X}^{*})=\{\bm{f}(\bm{x})\in\mathbb{R}^{m}|\bm{x}\in\bm{X}^{*}\}.$ (3)
The weakly Pareto set $\bm{X}_{w}^{*}$ and front $\bm{f}(\bm{X}_{w}^{*})$ can be defined accordingly. It is clear that $\bm{X}^{*}\subseteq\bm{X}_{w}^{*}$. Under mild conditions, the Pareto set $\bm{X}^{*}$ and the Pareto front $\bm{f}(\bm{X}^{*})$ are both $(m-1)$-dimensional manifolds in the decision space $\mathbb{R}^{n}$ and objective space $\mathbb{R}^{m}$, respectively (Hillermeier, 2001). Many multi-objective optimization methods have been proposed to find a single or a finite set of solutions in the Pareto set $\bm{X}^{*}$ for the problem (1). The scalarization approach and the adaptive gradient algorithm are two popular methods when all objective functions are differentiable.
Scalarization Approaches
Scalarization is a classical and popular method for tackling multi-objective optimization (Miettinen, 1999; Zhang & Li, 2007). The most straightforward approach is the simple linear scalarization (Geoffrion, 1967):
$\min_{\bm{x}\in\mathcal{X}}g^{(\text{LS})}(\bm{x}|\bm{\lambda})=\min_{\bm{x}\in\mathcal{X}}\sum_{i=1}^{m}\lambda_{i}f_{i}(\bm{x}),$ (4)
where $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{m})$ is a preference vector on the simplex $\bm{\Delta}^{m-1}=\{\bm{\lambda}|\sum_{i=1}^{m}\lambda_{i}=1\text{ and }\lambda_{i}\geq 0,\forall i\}$ over the $m$ objectives. Once a preference $\bm{\lambda}$ is given, we can obtain a solution by solving the single-objective scalarization problem (4). However, from the perspective of multi-objective optimization, this method can only find solutions on the convex hull of the Pareto front (Boyd & Vandenberghe, 2004; Ehrgott, 2005), and all solutions on the non-convex parts of the Pareto front will be missing (Das & Dennis, 1997). In fact, there is no guarantee that a general scalarization method can find all Pareto solutions, where the Tchebycheff scalarization method is an exception (Miettinen, 1999).
Tchebycheff Scalarization
In this work, we focus on the Tchebycheff (TCH) scalarization with promising theoretical properties (Bowman, 1976; Steuer & Choo, 1983):
$\min_{\bm{x}\in\mathcal{X}}g^{(\text{TCH})}(\bm{x}|\bm{\lambda})=\min_{\bm{x}\in\mathcal{X}}\max_{1\leq i\leq m}\left\{\lambda_{i}(f_{i}(\bm{x})-z^{*}_{i})\right\},$ (5)
where $\bm{\lambda}\in\bm{\Delta}^{m-1}$ is the preference vector and $\bm{z}^{*}\in\mathbb{R}^{m}$ is the ideal objective values (e.g., $\bm{z}^{*}_{i}=\min f_{i}(\bm{x})-\epsilon$ with a small $\epsilon>0$). Under mild conditions, its optimal solution $\bm{x}_{\bm{\lambda}}^{*}$ has the desirable pattern (Ehrgott, 2005):
$\lambda_{i}(f_{i}(\bm{x}_{\bm{\lambda}}^{*})-z^{*}_{i})=\lambda_{j}(f_{j}(\bm{x}_{\bm{\lambda}}^{*})-z^{*}_{j}),\quad\forall 1\leq i,j\leq m$ (6)
which naturally satisfies the exact preferred trad-off requirement (Mahapatra & Rajan, 2020). There is also a promising necessary and sufficient condition for TCH scalarization to find all (weakly) Pareto solutions (Choo & Atkins, 1983):
Theorem 2.3.
A feasible solution $\bm{x}\in\mathcal{X}$ is weakly Pareto optimal for the original problem (1) if and only if there exists a valid preference vector $\bm{\lambda}$ such that $\bm{x}$ is an optimal solution of the Tchebycheff scalarization problem (5).
In addition, if the optimal solution is unique for a given $\bm{\lambda}$, it is Pareto optimal (Miettinen, 1999). Although this scalarization approach is well known in the multi-objective optimization community for many years, it is rarely used for gradient-based optimization due to its nonsmoothness. Even when all objective functions $\left\{f_{1},\cdots,f_{m}\right\}$ are differentiable, the Tchebycheff scalarization (5) is non-differentiable, and hence suffers from a slow convergence rate by subgradient descent (Goffin, 1977) as illustrated in Figure 2. A detailed analysis will be given in the next section.
Adaptive Gradient Algorithms
Due to the shortcomings of scalarization methods, many adaptive gradient algorithms have been proposed to find a Pareto stationary solution (Fliege & Svaiter, 2000; Schäffler et al., 2002; Désidéri, 2012). We have the following definition:
Definition 2.4 (Pareto Stationary Solution).
A solution $\bm{x}\in\mathcal{X}$ is Pareto stationary if there exists a set of weights $\bm{\alpha}\in\bm{\Delta}^{m-1}=\{\bm{\alpha}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0\ \forall i\}$ such that the convex combination of gradients $\sum_{i=1}^{m}\alpha_{i}\nabla f_{i}(\bm{x})=\bm{0}$.
The Pareto stationarity is a necessary condition for Pareto optimality. In addition, when all objectives are convex and $\alpha_{i}>0\ \forall i$, it is also the Karush-Kuhn-Tucker (KKT) sufficient condition for Pareto optimality (Miettinen, 1999).
Adaptive gradient algorithms typically aim to find a valid gradient direction to improve the performance of all objectives simultaneously at each iteration. For example, the multiple gradient descent algorithm (MGDA) (Désidéri, 2012) calculates a gradient direction $\bm{d}_{t}=\sum_{i=1}^{m}\alpha_{i}\nabla f_{i}(\bm{x})$ at iteration $t$ that is a weighted convex combination of all objective gradients by solving the following problem:
$\min_{\alpha_{i}}||\sum\nolimits_{i=1}^{m}\alpha_{i}\nabla f_{i}(\bm{x}_{t})||_{2}^{2},$
$s.t.~{}\sum\nolimits_{i=1}^{m}\alpha_{i}=1,~{}~{}~{}\alpha_{i}\geq 0,\forall i=1,...,m.$ (7)
According to Désidéri (2012), the obtained $\bm{d}_{t}$ should either be a valid gradient direction to improve all objectives or $\bm{d}_{t}=\bm{0}$ which means the solution $\bm{x}_{t}$ is Pareto stationary. Therefore, we can use a simple gradient-based method to find a Pareto stationary solution (Fliege et al., 2019). In recent years, many work have adopted these gradient-based algorithms to tackle multi-objective optimization problems in the machine learning community (Sener & Koltun, 2018; Lin et al., 2019; Mahapatra & Rajan, 2020; Ma et al., 2020; Momma et al., 2022). Similar adaptive gradient methods are also developed to solve multi-task learning problems (Yu et al., 2020; Liu et al., 2021a, b, 2022; Navon et al., 2022; Senushkin et al., 2023; Lin et al., 2023a). Stochastic multi-objective optimization is another important topic for real-world applications (Liu & Vicente, 2021; Zhou et al., 2022; Fernando et al., 2023; Chen et al., 2023; Xiao et al., 2023).
However, at each iteration, these adaptive gradient algorithms typically need to calculate the gradients for each objective separately (e.g., via total $m$ different backpropagation) and solve a quadratic programming problem such as (7). This high pre-iteration complexity will lead to high computational overhead for solving large-scale problems such as training a deep neural network model. In addition, some recent works show that many adaptive gradient methods cannot significantly outperform a well-tuned or even random linear scalarization for deep multi-task learning problems (Kurin et al., 2022; Xin et al., 2022; Lin et al., 2022a; Royer et al., 2023), although linear scalarization fails to fully explore the whole Pareto front (Hu et al., 2023).
3 Smooth Tchebycheff Scalarization
Instead of proposing another adaptive gradient algorithm, this work revisits the straightforward scalarization approach. By leveraging the smooth optimization approach, we propose a lightweight and efficient smooth Tchebycheff scalarization with a fast convergence rate, low pre-iteration complexity, and promising theoretical properties for multi-objective optimization.
3.1 Smoothing Method for Nonsmooth Problem
We first analyze the nonsmoothness of classical Tchebycheff scalarization, and then briefly introduce the smooth optimization method that can be used to tackle this issue. We have the following definition of smoothness:
Definition 3.1 (Smoothness).
A function $g(\bm{x})$ is $L$-smooth if it has $L$-Lipschitz continuous gradient $\nabla g(\bm{x})$:
$||\nabla g(\bm{x})-\nabla g(\bm{y})||\leq L||\bm{x}-\bm{y}||,\forall\bm{x},\bm{y}\in\mathcal{X}.$ (8)
Nonsmoothness of Tchebycheff Scalarization
The nonsmoothness of classical Tchebycheff scalarization (5) comes from the non-differentiable $\mathtt{max}$ operator on all objectives. Even when all objective functions $f_{i}(\bm{x})$ are convex and smooth, the max function $g(x)=\max_{i}\{f_{i}(\bm{x})\}$ is convex but not smooth (Boyd & Vandenberghe, 2004). In other words, this function is not differentiable and only has subgradients with respect to the decision variable $\bm{x}$. It is well known that the subgradient method needs a large number of iterations in the order of $O(\frac{1}{\epsilon^{2}})$ to achieve an $\epsilon$-optimal solution ^$\hat{}\bm{x}$ (e.g., $||f(\hat{}\bm{x})-f(\bm{x}^{*})||\leq\epsilon$) for nonsmooth convex optimization (Goffin, 1977). In contrast, the required gradient descent iteration is $O(\frac{1}{\epsilon})$ for smooth convex optimization. The slow convergence rate of the Tchebycheff scalarization is also illustrated in Figure 2.
Smooth Optimization
Smooth optimization is a principled and powerful approach to handling nonsmooth optimization problems (Nesterov, 2005; Beck & Teboulle, 2012; Chen, 2012). According to Nesterov (2005), the large $O(\frac{1}{\epsilon^{2}})$ iterations is the worst-case estimate for a general nonsmooth problem, which can be significantly reduced if we have prior knowledge about the problem structure. According to Beck & Teboulle (2012) and Chen (2012), we have the following definition of smoothing function:
Definition 3.2 (Smoothing Function).
We call $g_{\mu}:\mathcal{X}\rightarrow\mathbb{R}$ a smoothing function of a continuous function $g$, if for any $\mu>0$, $g_{\mu}$ is continuously differentiable in $\mathcal{X}$ and satisfies the following conditions:
  1. 1.
    $\lim_{\bm{z}\rightarrow\bm{x},\mu\downarrow 0}g_{\mu}(\bm{z})=g(\bm{x}),\quad\forall\bm{x}\in\mathcal{X}$;
  2. 2.
    (Lipschitz smoothness with respect of $\bm{x}$) there exist constants $K$ and $\alpha>0$ irrelevant to $\mu$ such that $g_{\mu}$ is smooth on $\mathcal{X}$ with Lipschitz constant $L_{g_{\mu}}=K+\frac{\alpha}{\mu}$.
With a properly chosen smoothing parameter $\mu$, we can find an $\epsilon$-optimal solution to the original nonsmooth function $g(\bm{x})$ by optimizing the smoothing function $g_{\mu}(\bm{x})$ within $O(\frac{1}{\epsilon})$ iterations (Beck & Teboulle, 2012).
3.2 Smooth Tchebycheff Scalarization
By levering the infimal convolution smoothing method developed in Beck & Teboulle (2012), we propose the following smooth Tchebycheff (STCH) scalarization for multi-objective optimization:
$g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})=\mu\log\left(\sum_{i=1}^{m}e^{\frac{\lambda_{i}(f_{i}(\bm{x})-z^{*}_{i})}{\mu}}\right),$ (9)
where $\mu$ is the smoothing parameter, $m$ is the number of objectives, $\bm{\lambda}\in\bm{\Delta}^{m-1}$ and $\bm{z}^{*}\in\mathbb{R}^{m}$ are the preference vector and ideal objective values respectively as in the classical Tchebycheff scalarization (5). Like other scalarization methods, once a specific preference $\bm{\lambda}$ is given, we can directly optimize the STCH scalarization with a straightforward gradient descent algorithm as shown in Algorithm 1. Its pre-iteration complexity is much lower than the adaptive gradient methods that need to solve a quadratic programming problem such as (7) at each iteration.
The STCH scalarization (9) has many good theoretical properties. In the rest of this subsection, we illustrate why it is a promising alternative to the classical nonsmooth TCH scalarization (5). A theoretical analysis and discussion from the viewpoint of multi-objective optimization will be provided in the next subsection.
[…]
Figure 3: Smoothing a Nonsmooth Function: (a) The simple TCH scalarization function $g(x)=\max(f_{1}(x)=-x,f_{2}(x)=x)$ and its corresponding STCH scalarization with different smoothing parameters $\mu=1$ and $\mu=0.2$. The $g(x)$ is tightly bounded from above and below with a small $\mu$. (b) The gradient and smoothed gradients. TCH scalarization does not have gradient at $x=0$ while STCH is differentiable everywhere.
First of all, the STCH scalarization is a proper smooth approximation to the classical nonsmooth TCH counterpart:
Proposition 3.3 (Smooth Approximation).
The smooth Tchebycheff scalarization $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$ is a smoothing function of the classical Tchebycheff scalarization $g^{(\text{TCH})}(\bm{x}|\bm{\lambda})$ that satisfies Definition 3.2.
In other words, it shares similar properties with TCH scalarization that is promising for multi-objective optimization, while it is much easier to optimize via gradient-based methods due to its differentiable nature. This proposition is a direct corollary of Proposition 4.1 in Beck & Teboulle (2012). It is easy to check that the classic TCH scalarization $g^{(\text{TCH})}(\bm{x}|\bm{\lambda})$ is properly bounded from above and below:
Proposition 3.4 (Bounded Approximation).
For any solution $\bm{x}\in\mathcal{X}$, we have
$g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})-\mu\log{m}\leq g^{(\text{TCH})}(\bm{x}|\bm{\lambda})\leq g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda}).$ (10)
This means that the STCH scalarization $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$ is a uniform smooth approximation (Nesterov, 2005) of $g^{(\text{TCH})}(\bm{x}|\bm{\lambda})$. When the smoothing parameter $\mu$ is small enough, the point-wise bound is tight for all $\bm{x}\in\mathcal{X}$. This is also the case for the optimal solution $\bm{x}^{*}=\operatorname*{arg\,min}g^{(\text{TCH})}(\bm{x}|\bm{\lambda})$. An illustration can be found in Figure 3. This good property motivates us to find an approximate optimal solution for the nonsmooth $g^{(\text{TCH})}(\bm{x}|\bm{\lambda})$ by optimizing the smooth counterpart $g^{(\text{STCH})}(\bm{x}|\bm{\lambda})$. We have the following proposition:
Lemma 3.5 (Convexity).
If all objective functions $\{f_{1}(\bm{x}),\ldots,f_{m}(\bm{x})\}$ are convex, then the STCH scalarization $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$ with any valid $\mu$ and $\bm{\lambda}$ is convex.
Proposition 3.6 (Iteration Complexity).
If all objective functions are convex, with a properly chosen smoothing parameter $\mu$, we can obtain an $\epsilon$-optimal solution to the nonsmooth $g^{(\text{TCH})}_{\mu}(\bm{x}|\bm{\lambda})$ within $O(\frac{1}{\epsilon})$ iterations by solving the smooth counterpart $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$.
This proposition can be established by combining the above results on smooth approximation (Proposition 3.3), bounded approximation guarantee (Proposition 3.4) and the convexity of STCH scalarization (Lemma 3.5), as well as the convergence rate of fast gradient-based methods for smooth convex optimization in a straightforward manner. A seminal proof can be found in Theorem 3.1 of Beck & Teboulle (2012) for general smooth optimization. This required number of iterations is much lower than $O(\frac{1}{\epsilon^{2}})$ for directly optimizing the nonsmooth TCH scalarization (5) with subgradient descent.
Algorithm 1 STCH for Multi-Objective Optimization
1:  Input: Preference $\bm{\lambda}$, Initial $\bm{x}_{0}$, Step Size $\{\eta_{t}\}_{t=0}^{T}$
2:  for $t=1$ to $T$ do
3:     $\bm{x}_{t}=\bm{x}_{t-1}-\eta_{t}\nabla g^{(\text{STCH})}_{\mu}(\bm{x}_{t-1}|\bm{\lambda})$
4:  end for
5:  Output: Final Solution $\bm{x}_{T}$
3.3 Properties for Multi-Objective Optimization
In addition to being a promising smooth approximation of classical Tchebycheff scalarization, the smooth Tchebycheff scalarization itself also has several good theoretical properties for multi-objective optimization. First of all, its optimal solution is (weakly) Pareto optimal:
Theorem 3.7 (Pareto Optimality of the Solution).
The optimal solution of STCH scalarization (9) is weakly Pareto optimal of the original multi-objective optimization problem (1). In addition, the solution is Pareto optimal if either
  1. 1.
    all preference coefficients are positive ($\lambda_{i}>0\ \forall i$) or
  2. 2.
    the optimal solution is unique.
Proof Sketch.
These properties can be proved by contradiction based on Definition 2.2 for (weakly) Pareto optimality and the form of STCH Scalarization (9). A detailed proof can be found in Appendix A.1. ∎
According to Theorem 3.7, for any smoothing parameter $\mu$, the solution of STCH scalarization (9) with any valid preference $\bm{\lambda}$ is (weakly) Pareto optimal. However, to find all Pareto solutions, we have an additional requirement on the smoothing parameter $\mu$:
Theorem 3.8 (Ability to Find All Pareto Solutions).
Under mild conditions, there exists a $\mu^{*}$ such that, for any $0<\mu<\mu^{*}$, every Pareto solution of the original multi-objective optimization problem (1) is an optimal solution of the STCH scalarization problem (9) with some valid preference $\bm{\lambda}$.
Proof Sketch.
We can prove this theorem by analyzing the geometry relation between the level surface of STCH scalarization and the Pareto front. The key step is to find a theoretical upper bound of $\mu^{*}$ such that the level surface of STCH scalarization with any $\mu<\mu^{*}$ is totally enclosed by the Pareto front. The mild conditions are formally defined in Assumption A.1, and detailed proof can be found in Appendix A.2. ∎
Theorem 3.8 provides a sufficient condition for STCH scalarization to find all Pareto solutions of the multi-objective optimization problem (1). It can be treated as a smooth version of Theorem 2.3 for classical TCH scalarization. We have a stronger corollary for the convex Pareto front:
Corollary 3.9.
If the Pareto front is convex, then for any $\mu$, every Pareto solution of the original multi-objective optimization problem (1) is an optimal solution of the STCH scalarization problem (9) with some valid preference $\bm{\lambda}$.
The above analyses are for the global optimal properties of STCH scalarization. However, many real-world multi-objective optimization problems could be complicated and highly non-convex. In this case, we have the following local convergence guarantee.
Theorem 3.10 (Convergence to Pareto Stationary Solution).
If there exists a solution ^$\hat{}\bm{x}$ such that $\nabla g^{(\text{STCH})}_{\mu}(\hat{}\bm{x}|\bm{\lambda})=0$, then ^$\hat{}\bm{x}$ is a Pareto stationary solution of the original multi-objective optimization problem (1).
Proof Sketch.
This theorem can be proved by analyzing the form of the gradient for STCH scalarization. A crucial step is to show that the situation of Pareto stationarity in Definition 2.4 is satisfied when $\nabla g^{(\text{STCH})}_{\mu}(\bm{x}|\hat{}\bm{\lambda})=0$. A detailed proof can be found in Appendix A.3. ∎
There is a rich set of efficient gradient-based methods (Nocedal & Wright, 1999) that can be used to obtain a stationary solution for $g^{\text{(STCH)}}_{\mu}(\bm{x}|\bm{\lambda})$. Therefore, our proposed gradient-based STCH scalarization approach has a Pareto stationary guarantee similar to those of adaptive gradient algorithms, while allowing decision-makers to assign their preference among the objectives easily. Additionally, at each iteration, STCH scalarization does not need to calculate separate gradients for each objective and can also avoid solving a quadratic programming problem such as (7).
3.4 STCH Scalarization for Pareto Set Learning
Classic scalarization and adaptive gradient methods all focus on finding a single or finite set of Pareto solutions, but the whole Pareto set could be an $(m-1)$-dimensional manifold that contains infinite solutions (Hillermeier, 2001). Recently, some works have been proposed to build a model to approximate the Pareto set or front (Parisi et al., 2014; Dosovitskiy & Djolonga, 2019; Lin et al., 2020b; Navon et al., 2021; Ruchte & Grabocka, 2021; Lin et al., 2022b, c; Chen & Kwok, 2022; Jain et al., 2023; Lin et al., 2023b; Zhang et al., 2023). A Pareto set model can be defined as:
$\bm{x}^{*}(\bm{\lambda})=h_{\bm{\theta}}(\bm{\lambda})=\operatorname*{arg\,min}_{\bm{x}\in\mathcal{X}}g(\bm{x}|\bm{\lambda}),\forall\bm{\lambda}\in\Delta^{{m-1}},$ (11)
where a model $h_{\bm{\theta}}(\bm{\lambda})$ with parameter $\bm{\theta}$ maps any valid preference $\bm{\lambda}\in\Delta^{{m-1}}$ to its corresponding optimal solution $\bm{x}^{*}(\bm{\lambda})$ with respect to some scalarization function $g(\bm{x}|\bm{\lambda})$. The quality of this set model heavily depends on the scalarization function it uses. For example, a set model with linear scalarization might miss the solutions on the non-convex part of the Pareto front, and the classical Tchebycheff scalarization will lead to nonsmooth optimization for set model learning. Our proposed STCH scalarization $g^{\text{(STCH)}}_{\mu}(\bm{x}|\bm{\lambda})$ can serve as a promising drop-in replacement to efficiently learn a better Pareto set model.
Table 2: Results on the NYUv2 dataset.
Segmentation Depth Estimation Surface Normal Prediction $\Delta_{p}\uparrow$
[…]
Single Task Baseline
[…]
Adaptive Gradient Method
[…]
GradNorm 20.09 52.06 0.7200 0.2800 24.83 18.86 30.81 57.94 69.73 -11.7
PCGrad 38.06 64.64 0.5550 0.2325 27.41 22.80 23.86 49.83 63.14 +1.11
GradDrop 39.39 65.12 0.5455 0.2279 27.48 22.96 23.38 49.44 62.87 +2.07
GradVac 37.53 64.35 0.5600 0.2400 27.66 23.38 22.83 48.66 62.21 -0.49
IMTL-G 39.35 65.60 0.5426 0.2256 26.02 21.19 26.20 53.13 66.24 +4.77
CAGrad 39.18 64.97 0.5379 0.2229 25.42 20.47 27.37 54.73 67.73 +5.81
MTAdam 39.44 65.73 0.5326 0.2211 27.53 22.70 24.04 49.61 62.69 +3.21
Nash-MTL 40.13 65.93 0.5261 0.2171 25.26 20.08 28.40 55.47 68.15 +7.65
MetaBalance 39.85 65.13 0.5445 0.2261 27.35 22.66 23.70 49.69 63.09 +2.67
[…]
Aligned-MTL 40.82 66.33 0.5300 0.2200 25.19 19.71 28.88 56.23 68.54 +8.16
IMTL 41.19 66.37 0.5323 0.2237 26.06 20.77 26.76 53.48 66.32 +6.45
[…]
Adaptive Loss Method
UW 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 +0.91
DWA 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 +1.93
IMTL-L 39.78 65.27 0.5408 0.2347 26.26 20.99 26.42 53.03 65.94 +4.39
IGBv2 38.03 64.29 0.5489 0.2301 26.94 22.04 24.77 50.91 64.12 +2.11
Scalarization Method
EW 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 +0.88
GLS 39.78 65.63 0.5318 0.2272 26.13 21.08 26.57 52.83 65.78 +5.15
RLW 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 -2.16
TCH 34.09 59.76 0.5931 0.2524 28.30 23.99 21.87 47.13 60.61 -5.67
[…]
4 Experimental Studies
We evaluate the proposed STCH scalarization approach on finding a single Pareto solution and the whole Pareto set.
4.1 Multi-Task Learning
The STCH scalarization can be used to find a solution with balanced trade-off for multi-task learning problems.
Baseline Methods
We compare the STCH scalarization with the single-task learning (STL) baseline and (1) other scalarization methods: equal weight (EW), geometric mean loss (GLS) (Chennupati et al., 2019), random weights (RLW) (Lin et al., 2022a), and classical Tchebycheff scalarization (TCH); (2) adaptive loss methods: UW (Kendall et al., 2018), DWA (Liu et al., 2019), IMTL-L, and IGBv2 (Dai et al., 2023); (3) adaptive gradient methods: MGDA (Sener & Koltun, 2018), GradNorm (Chen et al., 2018), PCGrad (Yu et al., 2020), GradDrop (Chen et al., 2020), GradVac (Wang & Tsvetkov, 2021), IMTL-G, CAGrad (Liu et al., 2021a), MTAdam (Malkiel & Wolf, 2021), Nash-MTL (Navon et al., 2022), MetaBalance (He et al., 2022), MoCo (Fernando et al., 2023), Alighed-MTL (Senushkin et al., 2023), and DB-MTL (Lin et al., 2023a); (4) hybrid loss-gradient balancing method: IMTL (Liu et al., 2021b). All methods are implemented with the LibMTL library (Lin & Zhang, 2023). The problem and model training details can be found in Appendix D.1.
Evaluation Metrics
In addition to reporting the commonly used metrics for each task, we also report the relative average performance improvement (Maninis et al., 2019; Vandenhende et al., 2021) of each method over the STL baseline $\Delta_{p}=\frac{1}{T}\sum_{i}^{T}\Delta_{p,t}$ with $\Delta_{p,t}=100\%\times\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}(-1)^{s_{t,i}}\frac{M_{t,i}-M^{\text{STL}}_{t,i}}{M^{\text{STL}}_{t,i}}$, where $T$ is the number of tasks, $N_{t}$ is the number of metrics for task $t$, $M_{t,i}$ is the $i$-th metric of the MTL method on task $t$, and $s_{t,i}=0$ if a large value is better for the $i$-th metric on task $t$ and $1$ otherwise. A simple uniform preference is used for STCH. All methods are evaluated on three popular MTL datasets:
NYUv2 (Silberman et al., 2012) is an indoor scene understanding dataset with $3$ tasks on semantic segmentation, depth estimation, and surface normal prediction. The result in Table 2 shows that STCH scalarization can find a good trade-off solution with the second best overall $\Delta_{p}$ among all methods, which is only outperformed by a very recently proposed adaptive gradient method DB-MTL. Its large performance improvement over the classic TCH scalarization counterpart fully demonstrates the importance of smoothness for multi-objective optimization.
Office-31 (Saenko et al., 2010) is an image classification dataset across $3$ domains (Amazon, DSLR, and Webcam). According to Table 4, STCH scalarization can achieve the best overall performance among all methods. In fact, it is the only method that can dominate the STL baseline on all tasks. The full results can be found in Appendix E.1.
QM9 (Ramakrishnan et al., 2014) is a molecular property prediction dataset with 11 tasks. The results in Table 4 suggest that STCH scalarization has a very-close-to-best overall $\Delta_{p}$ with DB-MTL, while enjoying a much faster run time. The full results can be found in Appendix E.2.
Given its simplicity and promising performance, STCH scalarization can serve as a strong baseline for MTL.
[…]
Figure 4: The learned Pareto fronts for the 3-objective rocket injector design problem with different scalarization methods.
Table 3: Results (hypervolume difference $\Delta\text{HV}\downarrow$) on $6$ synthetic benchmark problems and $5$ real-world engineering design problems.
F1 F2 F3 F4 F5 F6 BarTruss HatchCover DiskBrake GearTrain RocketInjector
LS 1.64e-02 1.37e-02 9.40e-02 2.26e-01 1.72e-01 2.54e-01 8.03e-03 7.89e-03 4.05e-02 4.01e-03 1.42e-01
[…]
Table 4: Results on the Office-31 and QM9 datasets.
Office-31 QM9
$\Delta_{p}\uparrow$ $\Delta_{p}\uparrow$ $T\downarrow$
Single Task Baseline
STL 0.00 0.00 -
Adaptive Gradient Method
MGDA -0.27$\pm$0.15 -103.0$\pm$8.62 5.70
GradNorm -0.59$\pm$0.94 -227.5$\pm$1.85 4.47
PCGrad -0.68$\pm$0.57 -117.8$\pm$3.97 4.92
GradDrop -0.59$\pm$0.46 -191.4$\pm$9.62 1.34
GradVac -0.58$\pm$0.78 -150.7$\pm$7.41 5.06
IMTL-G -0.97$\pm$0.95 -1250$\pm$90.9 4.18
CAGrad -1.14$\pm$0.85 -87.25$\pm$1.51 4.71
MTAdam -0.60$\pm$0.93 -1403$\pm$203 -
Nash-MTL +0.24$\pm$0.89 -73.92$\pm$2.12 6.53
MetaBalance -0.63$\pm$0.30 -125.1$\pm$7.98 -
MoCo +0.89$\pm$0.26 -1314$\pm$65.2 4.51
Aligned-MTL -0.90$\pm$0.48 -80.58$\pm$4.18 4.45
IMTL -1.02$\pm$0.92 -104.3$\pm$11.7 4.62
[…]
Adaptive Loss Method
UW -0.56$\pm$0.90 -92.35$\pm$13.9 1.00
DWA -0.70$\pm$0.62 -160.9$\pm$16.7 1.00
IMTL-L -0.63$\pm$0.58 -77.06$\pm$11.1 1.00
IGBv2 +0.56$\pm$0.25 -99.86$\pm$10.4 1.00
Scalarization Method
EW -0.61$\pm$0.67 -146.3$\pm$ 7.86 1.00
GLS -1.63$\pm$0.61 -81.16$\pm$15.5 1.00
RLW -0.59$\pm$0.95 -200.9$\pm$13.4 1.00
TCH -0.71$\pm$0.56 -252.2$\pm$16.6 1.00
[…]
4.2 Efficient Pareto Set Learning
The STCH scalarization can also be used to improve the performance of Pareto set learning.
Baseline Methods
We build and compare Pareto set learning models with different scalarization methods: (1) Linear Scalarization (LS), (2) Conditioned One-shot Multi-Objective Search (COSMOS) (Ruchte & Grabocka, 2021), (3) Exact Preference Optimization (EPO) (Mahapatra & Rajan, 2020), (4) classic TCH scalarization, and (5) our proposed STCH scalarization.
Evaluation Metric
Following the related work, we use the hypervolume difference ($\Delta\text{HV}$(Zitzler et al., 2007) to measure the qualities of approximate Pareto fronts learned by different methods. A lower $\Delta\text{HV}$ means that the learned Pareto front is closer to the ground truth Pareto front and hence has better performance. A detailed definition of $\Delta\text{HV}$ can be found in Appendix D.2.3.
Results and Analysis
We evaluate all methods on six synthetic benchmark problems (F1-F6) as well as five real-world engineering design problems for Bar Truss, Hatch Cover, Disk Brake, Gear Train, and Rocket Injector. According to Table 3, the Pareto set learning model with STCH scalarization performs the best on $10$ out of $11$ problems. The approximate Pareto fronts learned by different scalarization methods for the rocket injector design are shown in Figure 4. The Pareto front learned by STCH can find better trade-offs than others. These results demonstrate that STCH is a promising scalarization approach for Pareto set learning. Problem details can be found in Appendix D.2.1 and D.2.2.
Table 5: The results of STCH with different smooth parameters $\mu$ on $5$ real-world engineering desgin problems.
BarTruss HatchCover DiskBrake GearTrain RocketInjector
TCH 9.05e-03 1.01e-02 3.78e-02 3.91e-03 2.73e-02
STCH($\mu=0.01$) 6.94e-03 8.47e-03 3.12e-02 3.79e-03 1.95e-02
STCH($\mu=0.1$) 6.85e-03 8.76e-03 2.79e-02 3.17e-03 2.00e-02
STCH($\mu=0.5$) 6.21e-03 7.81e-03 2.96e-02 3.36e-03 2.24e-02
STCH($\mu=1$) 6.95e-03 8.02e-03 3.32e-02 3.62e-03 3.24e-02
4.3 Effect of Different Smooth Levels
The parameter $\mu$ controls the smoothing level of the STCH scalarization, which acts as a hyperparameter in our proposed method. In this work, we do not aggressively tune $\mu$ for each problem. To investigate the effect of $\mu$, we conduct a new experiment to test STCH with different smooth levels $\mu$ for the multi-objective engineering design problem. As shown in Table 5, STCH scalarization with a small $\mu$ works reasonably well, but there is no single $\mu$ that can achieve the best performance for all problems.
The smooth parameter $\mu$ does have an impact on the convergence rate of the gradient-based algorithm. Roughly speaking, a larger $\mu$ can yield a lower iteration complexity for the smoothed function, while a small enough $\mu$ is required to obtain a good approximate solution for the original non-smooth problem. To better leverage this trade-off, a few homotopy methods (Allen-Zhu & Hazan, 2016; Xu et al., 2016) have been proposed to gradually decrease $\mu$ in a stage-wise manner for the Nesterov smoothing approach (Nesterov, 2005). This method can reduce the iteration complexity from $O(1/\epsilon)$ to $\tilde{O}(1/\epsilon^{1-\theta})$ where $\theta\in(0,1]$ represents the local sharpness around the optimal solution.
However, the homotopy schedule could be problem-specific and should be tuned for each new problem. We do not observe a clear performance improvement by using the homotopy method and, therefore, simply use a fixed $\mu$ in all our experiments. How to adaptively adjust the smooth parameter $\mu$ for each problem during the optimization process could be an interesting research direction in future work.
5 Local and Global Optimality Guarantee
Although Theorem 3.8 provides a promising sufficient condition for STCH scalarization to find all Pareto solutions, the simple gradient-based algorithm can only guarantee to find Pareto stationary solutions for a general multi-objective optimization problem as in Theorem 3.10. We would like to make the following remarks on this theoretical guarantee:
Similar Guarantee with the Other Methods
The global optimality guarantee for STCH is on par with the optimality guarantee for the classic TCH scalarization counterpart as in Theorem 2.3. Every Pareto optimal solution can be found by TCH scalarization, but the subgradient method can only guarantee to find a Pareto stationary solution. This guarantee is also on par with the adaptive gradient methods that are popular in the current literature, but our proposed method has a much lower pre-iteration complexity. Under a stronger assumption, such as all objectives are convex (and hence the STCH function is convex), the proposed method can find the global optimal solution.
Advanced Method for Global Optimality
It could be extremely hard, if not impossible, to provide a convergence guarantee to find the global optimal solution for a general non-convex optimization problem. Some advanced global optimization methods, such as homotopy/graduated optimization (Dunlavy & O’Leary, 2005; Hazan et al., 2016), could be helpful, but it is still a very difficult open problem for gradient-based optimization. We will explore this direction in future work. On the other hand, a local optimal solution is usually good enough for deep learning applications. According to the experimental results, our proposed method achieves promising performance on the MTL problems, while enjoying a significantly faster runtime than those adaptive gradient methods.
6 Conclusion, Limitation, and Future Work
Conclusion
In this work, we have proposed a novel smooth Tchebycheff (STCH) scalarization approach for efficient gradient-based multi-objective optimization. It has a fast convergence rate and low pre-iteration complexity, while enjoying many good theoretical properties. Experimental studies also show that STCH scalarization can obtain promising performance on various multi-objective optimization problems. We believe it can serve as a powerful yet straightforward method to tackle differentiable multi-objective optimization.
Limitation
This work only focuses on unconstrained and deterministic multi-objective optimization. It could be interesting to investigate how to properly handle different constraints and tackle stochastic optimization with STCH scalarization in future work. A brief discussion on these directions can be found in Appendix B.4 and B.5.
Acknowledgements
The work described in this paper was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (GRF Project No. CityU11215622), the National Natural Science Foundation of China (Grant No. 62106096), the Natural Science Foundation of Guangdong Province (Grant No. 2024A1515011759), the National Natural Science Foundation of Shenzhen (Grant No. JCYJ20220530113013031).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Detailed Proofs
A.1 Proof of Theorem 3.7
Theorem 3.7 (Pareto Optimality of the Solution). The optimal solution of STCH scalarization (9) is weakly Pareto optimal. In addition, the solution is Pareto optimal if either
  1. 1.
    all preference coefficients are positive ($\lambda_{i}>0\ \forall i$) or
  2. 2.
    the optimal solution is unique.
Proof.
Weakly Pareto Optimality: We first prove the optimal solution of STCH scalarization is weakly Pareto optimal by contradiction. Let $\bm{x}^{*}$ be an optimal solution for the STCH scalarization $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$ with valid preference $\bm{\lambda}$, we have:
$\bm{x}^{*}=\operatorname*{arg\,min}_{\bm{x}\in\mathcal{X}}\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\bm{x})-z^{*}_{i})}/{\mu}}\right).$ (12)
Suppose that $\bm{x}^{*}$ is not weakly Pareto optimal for the multi-objective optimization problem (1). According to Definition 2.2 on the (weakly) Pareto optimality, there exists a valid solution ^$\hat{}\bm{x}\in\mathcal{X}$ such that $\bm{f}(\hat{}\bm{x})\prec_{\text{strict}}\bm{f}(\bm{x}^{*})$. In other words, we have:
$f_{i}(\hat{}\bm{x})<f_{i}(\bm{x}^{*})\quad\forall i\in\{1,...,m\}.$ (13)
Based on the above inequalities, it is easy to see:
$\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\hat{}\bm{x})-z^{*}_{i})}/{\mu}}\right)<\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\bm{x}^{*})-z^{*}_{i})}/{\mu}}\right)$ (14)
which contradicts the optimality of $\bm{x}^{*}$ for the STCH scalarization (12). Therefore, $\bm{x}^{*}$ should be weakly Pareto optimal for the original multi-objective optimization problem.
Then we provide proof of the two sufficient conditions for $\bm{x}^{*}$ to be Pareto optimal in a similar way.
1. All Positive Preference Coefficients: Suppose that all preference coefficients $\bm{\lambda}_{i}$ are positive. If $\bm{x}^{*}$ is not Pareto optimal, there exists a valid solution ^$\hat{}\bm{x}\in\mathcal{X}$ such that $\bm{f}(\hat{}\bm{x})\prec\bm{f}(\bm{x}^{*})$. In other words, we have:
$f_{i}(\hat{}\bm{x})\leq f_{i}(\bm{x}^{*}),\forall i\in\{1,...,m\}\text{ and }f_{j}(\hat{}\bm{x})<f_{j}(\bm{x}^{*}),\exists j\in\{1,...,m\}.$ (15)
Based on the above inequalities and the condition $\bm{\lambda}_{i}>0\ \forall i\in\{1,...,m\}$, it is easy to see $\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\hat{}\bm{x})-z^{*}_{i})}/{\mu}}\right)<\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\bm{x}^{*})-z^{*}_{i})}/{\mu}}\right)$ which contradicts the STCH optimality of $\bm{x}^{*}$ in (12). Therefore, $\bm{x}^{*}$ should be Pareto optimal for the original multi-objective optimization problem.
2. Uniqueness of the Solution: Suppose $\bm{x}^{*}$ is a unique optimal solution for the STCH scalarization. If $\bm{x}^{*}$ is not Pareto optimal, there exists a valid solution ^$\hat{}\bm{x}\in\mathcal{X}$ that satisfies the inequalities in (15). It is easy to check (it is different from (14)):
$\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\hat{}\bm{x})-z^{*}_{i})}/{\mu}}\right){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\leq}\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\bm{x}^{*})-z^{*}_{i})}/{\mu}}\right).$ (16)
On the contrary, the condition of $\bm{x}^{*}$ being a unique optimal solution for STCH scalarization indicates that
$\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\hat{}\bm{x})-z^{*}_{i})}/{\mu}}\right){\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}>}\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\bm{x}^{*})-z^{*}_{i})}/{\mu}}\right).$ (17)
These two inequalities above are contradictory, and therefore the solution $\bm{x}^{*}$ should be Pareto optimal.
A.2 Proof of Theorem 3.8
[…]
Figure 5: The level surfaces of smooth Tchebycheff (STCH) scalarization and different Pareto fronts.
Following the work on convexification of the Pareto front (Li, 1996; Goh & Yang, 1998), we prove this theorem by analyzing the relationship between the level surface of the STCH scalarization and the Pareto front as illustrated in Figure 5.
First, we express the Pareto front for a multi-objective optimization problem as the surface $\phi(\bm{y})=0$ where $\bm{y}\in\mathbb{R}^{m}$ is a vector in the objective space. Then we make the following assumption following the previous work:
Assumption A.1 (Li (1996)).
For each point $\bm{y}$ on the Pareto front, we assume:
  1. 1.
    without loss of generality, $\nabla\phi(\bm{y})\succ 0$, which means $\frac{\partial\phi(\bm{y})}{\partial y_{i}}>0$ for all $i=1,\ldots,m$; and
  2. 2.
    all second-order derivatives of $\phi(\bm{y})$ are bounded.
For the first assumption, we can easily reverse the sign of $\phi(\bm{y})$ if $\nabla\phi(\bm{y})\prec 0$. This assumption also indicates that $\bm{f}(\bm{X}_{w}^{*})\setminus\bm{f}(\bm{X}^{*})=\emptyset$ which means no solution can be weakly Pareto optimal but not Pareto optimal in this problem. The second assumption means that the eigenvalues of the Hessian $\nabla^{2}\phi(\bm{y})$ are finite. With these two assumptions, we are ready to give a proof for the theorem.
Theorem 3.8 (Ability to Find All Pareto Solutions). Under mild conditions, there exists a $\mu^{*}$ such that, for any $0<\mu<\mu^{*}$, every Pareto solution of the original multi-objective optimization problem (1) is an optimal solution of the STCH scalarization problem (9) with some valid preference $\bm{\lambda}$.
Proof.
We denote the level surface of the STCH scalarization by $\psi(\bm{y})$ with:
$\psi(\bm{y})=\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(y_{i}-z^{*}_{i})}/{\mu}}\right)-c=0.$ (18)
For any valid Pareto solution $\bm{y}$ on the Pareto front, suppose that the STCH level surface $\psi(\bm{y})=0$ with some preference $\bm{\lambda}$ is tangential to the Pareto front $\phi(\bm{y})=0$. To prove this theorem, it is equivalent to show $\bm{y}$ must be an optimal solution of the STCH scalarization problem (9). In other words, the STCH level surface must be enclosed by the Pareto front toward the origin. Since these two surfaces are tangential at $\bm{y}$, the curvature of the STCH level surface $\psi(\bm{y})=0$ should be larger than the curvature of the Pareto front $\phi(\bm{y})=0$. Some illustrations can be found in Figure 5. In the following, we give a sufficient condition for this case with a small enough $\mu^{*}$.
For better exposition and without loss of generality, we further assume that the level surface of STCH scalarization with equal preference $\bm{\lambda}=(1,1,\ldots,1)$ and ideal point $\bm{z}^{*}=0$ is tangential to the Pareto front at some point $\bm{y}^{*}$. This case can be easily generalized to all valid preference $\bm{\lambda}$ and ideal point $\bm{z}^{*}$.
Since the two surfaces $\psi(\bm{y})=0$ and $\phi(\bm{y})=0$ are tangential at $\bm{y}^{*}$, we have:
$\psi(\bm{y}^{*})=\phi(\bm{y}^{*}),\quad\nabla\psi(\bm{y}^{*})=\nabla\phi(\bm{y}^{*}).$ (19)
According to Assumption A.1, the Pareto front $\phi(\bm{y})$ is a smooth function with respect to $\bm{y}$. It is also easy to check that the STCH level surface $\psi(\bm{y})$ is also smooth. Therefore, we can expand both $\phi(\bm{y})$ and $\psi(\bm{y})$ near $\bm{y}^{*}$ with their Taylor series up to the second term:
$\phi(\bm{y})=\phi(\bm{y}^{*})+\nabla\phi(\bm{y}^{*})(\bm{y}-\bm{y}^{*})+\frac{1}{2}(\bm{y}-\bm{y}^{*})^{T}\nabla^{2}\phi(\bm{y}^{*})(\bm{y}-\bm{y}^{*})+o(||\bm{y}-\bm{y}^{*}||^{2}),$ (20)
$\psi(\bm{y})=\psi(\bm{y}^{*})+\nabla\psi(\bm{y}^{*})(\bm{y}-\bm{y}^{*})+\frac{1}{2}(\bm{y}-\bm{y}^{*})^{T}\nabla^{2}\psi(\bm{y}^{*})(\bm{y}-\bm{y}^{*})+o(||\bm{y}-\bm{y}^{*}||^{2}).$ (21)
Taking them together with (19), we have:
$\psi(\bm{y})-\phi(\bm{y})=\frac{1}{2}(\bm{y}-\bm{y}^{*})^{T}(\nabla^{2}\psi(\bm{y}^{*})-\nabla^{2}\phi(\bm{y}^{*}))(\bm{y}-\bm{y}^{*})+o(||\bm{y}-\bm{y}^{*}||^{2}).$ (22)
The STCH level surface $\psi(\bm{y})=0$ is enclosed by the Pareto front $\phi(\bm{y})=0$ towards the origin if $\psi(\bm{y})-\phi(\bm{y})\geq 0$ for all $\bm{y}$ in the neighborhood of $\bm{y}^{*}$. According to Assumption A.1, the term $\nabla^{2}\phi(\bm{y}^{*})$ is bounded. Therefore, it is important to analyze the term $\nabla^{2}\psi(\bm{y}^{*})=\nabla^{2}g^{\text{(STCH)}}_{\mu}(\bm{y}^{*})$ which is the Hessian matrix of STCH scalarization with respect to the objective value $\bm{y}$.
For STCH scalarization with a equal preference and ideal point at $\bm{0}$, we have
$\nabla_{\bm{y}}g^{\text{(STCH)}}_{\mu}=\nabla_{\bm{y}}\mu\log\left(\sum_{i=1}^{m}e^{{y_{i}}/{\mu}}\right)=\frac{e^{{\bm{y}}/{\mu}}}{\sum_{i}e^{{y_{i}}/{\mu}}}\in\mathbb{R}^{m}.$ (23)
It is easy to show the STCH scalarization is twice differentiable with
$\nabla_{\bm{y}}^{2}g^{\text{(STCH)}}_{\mu}=\frac{1}{\mu}\left(\frac{1}{\mathbf{1}^{T}\bm{z}}\operatorname{diag}(\bm{z})-\frac{1}{(\mathbf{1}^{T}\bm{z})^{2}}\bm{z}\bm{z}^{T}\right),\text{where }\bm{z}=e^{\bm{y}}.$ (24)
For any $\bm{v}\in\mathbb{R}^{m}$, we have
$\bm{v}^{T}\nabla_{\bm{y}}^{2}g^{\text{(STCH)}}_{\mu}\bm{v}$ $=\bm{v}^{T}\left[\frac{1}{\mu}\left(\frac{1}{\mathbf{1}^{T}\bm{z}}\operatorname{diag}(\bm{z})-\frac{1}{(\mathbf{1}^{T}\bm{z})^{2}}\bm{z}\bm{z}^{T}\right)\right]\bm{v}$ (25)
$=\frac{(\sum_{i=1}^{m}\bm{z}_{i}\bm{v}^{2}_{i})(\sum_{i=1}^{m}\bm{z}_{i})-(\sum_{i=1}^{m}\bm{v}_{i}\bm{z}_{i})^{2}}{\mu(\sum_{i=1}^{m}\bm{z}_{i})^{2}}.$ (26)
According to the Cauchy-Schwarz inequality, we have $(\sum_{i=1}^{m}\bm{v}_{i}\bm{z}_{i})^{2}\leq(\sum_{i=1}^{m}\bm{z}_{i}\bm{v}^{2}_{i})(\sum_{i=1}^{m}\bm{z}_{i})$ and the two sides are equal if and only if $\bm{v}=\gamma\bm{I}_{m}$ with any $\gamma\in\mathbb{R}$. It means $\bm{v}^{T}\nabla_{\bm{y}}^{2}g^{\text{(STCH)}}_{\mu}\bm{v}\geq 0$ for all $\bm{v}\in\mathbb{R}^{m}$. In other words, $\nabla_{\bm{y}}^{2}g^{\text{(STCH)}}_{\mu}$ is positive semi-definite ($\nabla_{\bm{y}}^{2}g^{\text{(STCH)}}_{\mu}\succeq 0$). Now we are ready to analyze the cases for different Pareto fronts.
Convex Pareto Front
If the Pareto set is convex (here it means cone-convex), all the eigenvalues of its Hessian $\nabla^{2}\phi(\bm{y})$ are negative for any valid $\bm{y}$. Since $\nabla^{2}\psi(\bm{y}^{*})$ is positive semi-definite, it is easy to check $\psi(\bm{y})-\phi(\bm{y})>0$ for every $\bm{y}$, which is agnostic to the smoothing parameter $\mu$. Therefore, for any $\mu>0$, every Pareto solution of the original multi-objective optimization problem (1) is an optimal solution of the STCH scalarization problem (9) with some valid preference $\bm{\lambda}$ (Corollary 3.9).
General Pareto Front
Let $\bm{v}=(\bm{y}-\bm{y}^{*})$, we can combine (22) and (26) to have
$\psi(\bm{y})-\phi(\bm{y})$ $=\frac{1}{2}(\bm{y}-\bm{y}^{*})^{T}(\nabla^{2}\psi(\bm{y}^{*})-\nabla^{2}\phi(\bm{y}^{*}))(\bm{y}-\bm{y}^{*})+o(||\bm{y}-\bm{y}^{*}||^{2})$ (27)
$=\frac{(\sum_{i=1}^{m}\bm{z}_{i}\bm{v}^{2}_{i})(\sum_{i=1}^{m}\bm{z}_{i})-(\sum_{i=1}^{m}\bm{v}_{i}\bm{z}_{i})^{2}}{2\mu(\sum_{i=1}^{m}\bm{z}_{i})^{2}}-\frac{1}{2}\bm{v}^{T}\nabla^{2}\phi(\bm{y}^{*})\bm{v}+o(||\bm{y}-\bm{y}^{*}||^{2})$ (28)
$\geq\frac{(\sum_{i=1}^{m}\bm{z}_{i}\bm{v}^{2}_{i})(\sum_{i=1}^{m}\bm{z}_{i})-(\sum_{i=1}^{m}\bm{v}_{i}\bm{z}_{i})^{2}}{2\mu(\sum_{i=1}^{m}\bm{z}_{i})^{2}}-K+o(||\bm{y}-\bm{y}^{*}||^{2})$ (29)
where $K$ is some constant since $\nabla^{2}\phi(\bm{y}^{*})$ is bounded according to Assumption A.1. According to (26) and the Cauchy-Schwarz inequality, if $\bm{v}=(\bm{y}-\bm{y}^{*})\neq\gamma\bm{I}_{m}$ with any $\gamma\in\mathbb{R}$, the first term of (29) is always positive. Therefore, there exists a small enough $\mu^{*}$ such that
$\psi(\bm{y})-\phi(\bm{y})>0,\forall 0<\mu<\mu^{*}.$ (30)
Furthermore, since $\bm{y}^{*}$ is the optimal Pareto solution with a equal preference $(1,1,\ldots,1)$, we have $y_{i}^{*}=y_{j}^{*}\forall 0\leq i,j\leq m$. If $\bm{v}=(\bm{y}-\bm{y}^{*})=\gamma\bm{I}_{m}$, the point $\bm{y}$ should also satisfy $y_{i}=y_{j}\forall 0\leq i,j\leq m$. However, according to the definition of Pareto front and the property of STCH scalarization, if $\bm{y}\neq\bm{y}^{*}$, $\bm{y}$ cannot be on the Pareto front or the STCH level surface. In other words, such $\bm{y}$ is not a valid neighbor point on the surface $\psi(\bm{y})$ or $\phi(\bm{y})$.
Therefore, this proof is completed. ∎
Corollary 3.9. If the Pareto front is convex, then for any $\mu$, every Pareto solution of the original multi-objective optimization problem (1) is an optimal solution of the STCH scalarization problem (9) with some valid preference $\bm{\lambda}$.
A.3 Proof of Theorem 3.10
Theorem 3.10 (Convergence to Pareto Stationary Solution). If there exists a solution ^$\hat{}\bm{x}$ such that $\nabla g^{(\text{STCH})}_{\mu}(\hat{}\bm{x}|\bm{\lambda})=0$, then ^$\hat{}\bm{x}$ is a Pareto stationary solution of the original multi-objective optimization problem (1).
Proof.
Let $\bm{y}=\bm{\lambda}(\bm{f}(\bm{x})-\bm{z}^{*})$ be an $m$-dimensional vector, we can rewrite the STCH scalarization as:
$g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})=\mu\log\left(\sum_{i=1}^{m}e^{{\lambda_{i}(f_{i}(\bm{x})-z^{*}_{i})}/{\mu}}\right)=\mu\log\left(\sum_{i=1}^{m}e^{{y_{i}}/{\mu}}\right).$ (31)
Then take the gradient of $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$ with respect to $\bm{y}$, we have:
$\nabla_{\bm{y}}g^{\text{(STCH)}}_{\mu}(\bm{x}|\bm{\lambda})=\nabla_{\bm{y}}\mu\log\left(\sum_{i=1}^{m}e^{{y_{i}}/{\mu}}\right)=\frac{e^{{\bm{y}}/{\mu}}}{\sum_{i}e^{{y_{i}}/{\mu}}}\in\mathbb{R}^{m},$ (32)
Therefore, according to the chain rule, the gradient of $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$ with respect to the solution $\bm{x}$ is:
$\nabla_{\bm{x}}g^{\text{(STCH)}}_{\mu}(\bm{x}|\bm{\lambda})$ $=\nabla_{\bm{y}}g^{\text{(STCH)}}_{\mu}(\bm{x}|\bm{\lambda})\cdot\frac{\partial\bm{y}}{\partial\bm{x}}$
$=\sum_{i=1}^{m}\frac{e^{{y_{i}}/{\mu}}}{\sum_{i}e^{{y_{i}}/{\mu}}}\nabla\lambda_{i}(f_{i}(\bm{x})-z_{i}^{*})$ (33)
$=\sum_{i=1}^{m}\frac{\lambda_{i}e^{{y_{i}}/{\mu}}}{\sum_{i}e^{{y_{i}}/{\mu}}}\nabla f_{i}(\bm{x}).$ (34)
It is easy to check $w_{i}=\frac{\lambda_{i}e^{{y_{i}}/{\mu}}}{\sum_{i}e^{{y_{i}}/{\mu}}}\geq 0,\forall i$. Therefore, let $s=\sum_{i=1}^{m}w_{i}$ and $\bar{w}_{i}=w_{i}/s$, we have:
$\nabla_{\bm{x}}g^{\text{(STCH)}}_{\mu}(\bm{x}|\bm{\lambda})=\sum_{i=1}^{m}w_{i}\nabla f_{i}(\bm{x})=s\sum_{i=1}^{m}\bar{w}_{i}\nabla f_{i}(\bm{x}),$ (35)
where $s>0$ and ^$\hat{}\bm{w}\in\bm{\Delta}^{m-1}=\{\hat{}\bm{w}|\sum_{i=1}^{m}\bar{w}_{i}=1,\bar{w}_{i}\geq 0\ \forall i\}$. For a solution ^$\hat{}\bm{x}$, if $\nabla g^{(\text{STCH})}_{\mu}(\hat{}\bm{x}|\bm{\lambda})=0$, we have:
$\nabla_{\bm{x}}g^{\text{(STCH)}}_{\mu}(\hat{}\bm{x}|\bm{\lambda})=s\sum_{i=1}^{m}\bar{w}_{i}\nabla f_{i}(\hat{}\bm{x})=0$ (36)
$\Longrightarrow_{\text{divided by }s}\sum_{i=1}^{m}\bar{w}_{i}\nabla f_{i}(\hat{}\bm{x})=0.$
According to Definition 2.4 of Pareto stationarity, the solution ^$\hat{}\bm{x}$ is a Pareto stationary solution. ∎
Appendix B Practical Implementation for STCH Scalarization
B.1 Computational Stability
For the STCH scalarization, we have
$g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})=\mu\log\left(\sum_{i=1}^{m}e^{\frac{y_{i}}{\mu}}\right)$ (37)
$\nabla_{\bm{x}}g^{\text{(STCH)}}_{\mu}(\bm{x}|\bm{\lambda})=\sum_{i=1}^{m}\frac{\lambda_{i}e^{{y_{i}}/{\mu}}}{\sum_{i}e^{{y_{i}}/{\mu}}}\nabla f_{i}(\bm{x}),$ (38)
where $\bm{y}_{i}=\bm{\lambda}_{i}(\bm{f}_{i}(\bm{x})-z^{*}_{i})$. These two terms might both have computational stability issues with a very small smoothing parameter $\mu$.
Following the stable technique proposed in Nesterov (2005), we first calculate $\tilde{y}=\max_{\forall i}y_{i}$, and then set $\hat{y}_{i}=y_{i}-\tilde{y}$ for all $i$. Then we have a stabalized version of the STCH scalarization and its gradient:
$\hat{g}^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})=\mu\log\left(\sum_{i=1}^{m}e^{{\hat{y}_{i}}/{\mu}}\right)$ (39)
$\nabla_{\bm{x}}\hat{g}^{\text{(STCH)}}_{\mu}(\bm{x}|\bm{\lambda})=\sum_{i=1}^{m}\frac{\lambda_{i}e^{{\hat{y}_{i}}/{\mu}}}{\sum_{i}e^{{\hat{y}_{i}}/{\mu}}}\nabla f_{i}(\bm{x}).$ (40)
B.2 Objective Normalization
In many real-world application problems, the objective functions could be in very different scales. It could be very difficult for the user to assign a meaningful preference among the objectives. In this case, we can first normalize each objective
$\hat{f}_{i}=\frac{f_{i}-f_{\text{i, min}}}{f_{\text{i, max}}-f_{\text{i, min}}}$ (41)
with known or predicted lower bounds and upper bounds, and then apply STCH scalarization on the normalized objective ^$\hat{}\bm{f}(\bm{x})=(\hat{f}_{1}(\bm{x}),\ldots,f_{m}(\bm{x}))$. For all multi-task learning problems, following the setting from Dai et al. (2023), we set $f_{\text{i, max}}$ to be the training loss for the $i$-th task at the end of the second epoch and $f_{\text{i, min}}=0$.
B.3 Setting the Smoothing Parameter $\mu$
For the STCH scalarization, there is an remained challenge on how to set the smoothing parameter. According to Nesterov (2005) and Beck & Teboulle (2012), the smoothing parameter $\mu$ for smooth optimization should be set properly based on the properties of the original function (e.g., Lipschitz constant $L$), which is usually unknown. Some practical algorithms have been proposed to continually solve a smoothed surrogate sequence with a schedule of large to small smoothing parameters (Xu et al., 2016). However, this approach could lead to extra computational overhead. In this work, we simply use a fixed smoothing parameter in all experiments, and find that the fixed $\mu$ work pretty well.
B.4 Constrains Handling
This work mainly focuses on the unconstrained multiobjective optimization problem, while the constrained problem could also be important for many real-world problems. Extra approaches could be needed for STCH to handle the constraints.
A constrained multiobjective optimization problem can be defined as
$\min_{x\in\Omega}\ f(x)=(f_{1}(x),f_{2}(x),\cdots,f_{m}(x))$ (42)
with $\Omega=\{x\in R^{n}:g_{j}(x)\leq 0,j=1,\ldots,p,\quad h_{l}(x)=0,l=1,\dots,q\}$, where $\{g_{j}(x)\}$ and $\{h_{l}(x)\}$ are the inequality and equality constraints. For differentiable multiobjective optimization, we assume that all objectives $\{f_{i}(x)\}$ and constraints $\{g_{j}(x)\},\{h_{l}(x)\}$ are continuously differentiable.
Following previous work on gradient-based multiobjective optimization (Fliege & Vaz, 2016), we can reformulate the constrained multiobjective optimization problem into an unconstrained problem with additional objectives:
$\min_{x\in R^{n}}\ (f(x),g^{+}(x),|h(x)|)=(f_{1}(x),\cdots,f_{m}(x),g^{+}_{1}(x),\cdots,g^{+}_{p}(x),|h_{1}(x)|,\cdots,|h_{q}(x)|)$ (43)
where $g^{+}_{j}(x)=\max\{0,g_{j}(x)\}$. A feasible solution $x$ for the original constrained problem satisfies $\{g^{+}_{j}(x)=0,j=1,\ldots,p,\quad|h_{l}(x)|=0,l=1,\dots,q\}$. Instead of the original problem with $m$ objectives and $p+q$ constraints, we now have an unconstrained problem with $m+p+q$ objectives and 00 constraints. With our proposed STCH scalarization, we have the following problem:
$g^{(\text{STCH})}_{\mu}(x|\lambda,\sigma)=\mu\log\left(\sum_{i=1}^{m}e^{\frac{\lambda_{i}(f_{i}(x)-z^{*}_{i})}{\mu}}+\sum_{j=1}^{p}e^{\frac{\sigma g^{+}_{j}(x)}{\mu}}+\sum_{l=1}^{q}e^{\frac{\sigma|h_{l}(x)|}{\mu}}\right),$ (44)
where $\sigma\geq 0$ is a penalty parameter. We can also use the smooth counterparts for $g^{+}_{j}(x)$ and $|h_{l}(x)|$ in the STCH scalarization above to achieve a better convergence rate. In addition to this soft constraint handling approach, there could be many other possible constraint handling methods that work for STCH, which is an important research direction in the future.
B.5 Potential Extension for Stochastic Multi-Objective Optimization
In this work, we mainly focus on deterministic multi-objective optimization setting, and leave the study for stochastic optimization as an important future work. We would like to make the following remarks on this concern:
  • Stochastic multi-objective optimization itself is an important research topic that was not yet well explored in the past decades. Most existing methods only study the deterministic setting. Until recently, a few work have been proposed to investigate the stochastic adaptive gradient method (Liu & Vicente, 2021; Zhou et al., 2022; Fernando et al., 2023). In this work, instead of yet another adaptive gradient method, we focus on proposing a lightweight, efficient, and theoretically solid scalarization method.
  • Although developed from the deterministic setting, our proposed STCH scalarization method can outperform MOCO (Fernando et al., 2023), a state-of-the-art stochastic adaptive gradient method on all multi-task learning datasets (for performance and run time), which are essentially stochastic optimization problems. These results validate that our proposed method can be used to tackle stochastic optimization problems.
  • In addition, how to design an efficient algorithm for stochastic nonsmooth optimization problem is also an emerging research topic. A recent work (Wang & Zhang, 2023) investigates efficient accelerated gradient methods for the classic smoothing method proposed in Nesterov (2005) and Beck & Teboulle (2012). We might leverage this recent development to investigate efficient scalarization methods for stochastic multi-objective optimization.
Appendix C Discussion with Related Work on Minimax Optimization
The Tchebycheff scalarization approach is naturally a minimax optimization problem. In this section, we briefly discuss the relation between our proposed STCH method and the related algorithms for minimax optimization.
Problem Formulation
A general minimax optimization problem is in the form:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}f(x,y),$
where $\mathcal{X}\in R^{n}$ and $\mathcal{Y}\in R^{m}$ are both closed and bounded convex sets, while $f:\mathcal{X}\times\mathcal{Y}\rightarrow R$ is a smooth function. From the viewpoint of Tchebycheff scalarization, we are more interested in the problem with a specific linear structure on $y$:
$\min_{x\in\mathcal{X}}\max_{y\in\Delta^{m-1}}F(x)^{T}y,$
where $F(x)=(f_{1}(x),f_{2}(x),\cdots,f_{m}(x))^{T}$ and $\Delta^{m-1}=\{y|\sum_{i=1}^{m}y_{i}=1,y_{i}\geq 0\ \forall i\}$ is a probability simplex. It is easy to check that the optimal solution of the above problem is also optimal for the Tchebycheff scalarization problem $\min_{x\in\mathcal{X}}\max_{1\leq i\leq m}f_{i}(x)$ since the linear optimization problem over the simplex is always optimized at one of its vertices (Boyd & Vandenberghe, 2004). Here we drop the preference $\lambda$ and reference point $z^{*}$ for better exposition.
Algorithms for Minimax Optimization
There are many algorithms to solve the minimax optimization problems, such as some classic methods (Nemirovski, 2004; Nesterov, 2007; Monteiro & Svaiter, 2010). In recent years, different algorithms have been proposed to tackle the minimax problems in different settings (Palaniappan & Bach, 2016; Gidel et al., 2017; Mertikopoulos et al., 2019). Rather than those algorithms with a double-loop or even triple-loop structure, we are more interested in the lightweight single-loop algorithm for minimax optimization such as Gradient Descent Ascent (GDA) (Lin et al., 2020a). However, the simple GDA algorithm could oscillate during the optimization process (Mokhtari et al., 2020) and fail to converge even for a simple bilinear problem $\min_{x\in R^{n}}\max_{y\in R^{n}}x^{T}y$ (Letcher et al., 2019). Extra efforts, such as alternating gradient projection (Xu et al., 2023) or a smoothing method (Zhang et al., 2020), are needed for GDA.
Comparison and Discussion
The smoothed gradient descent ascent (smoothed-GDA) (Zhang et al., 2020) algorithm is the most related minmax optimization algorithm to our proposed STCH method. The key approach in smoothed-GDA is to add an extra smoothed regularization term to the objective function:
$K(x,z;y)=f(x,y)+\frac{p}{2}||x-z||^{2}.$ (45)
To optimize the above function, in addition to the original $x$ and $y$, smoothed-GDA maintains and updates an auxiliary sequence $\{z^{t}\}$ where $z^{t+1}=z^{t}+\beta(x^{t+1}-z^{t})$ throughout the optimization process. The terms $p>0$ and $\beta>0$ are hyperparameters, and the smoothed-GDA reduces to the standard GDA when $\beta=1$.
Compared to smoothed-GDA, our proposed STCH approach does not require maintaining an extra sequence $\{z^{t}\}$ which could lead to high memory cost (e.g., an additional deep neural network with millions of parameters). For Pareto set learning $x^{*}(\lambda)=h_{\theta}(\lambda)=argmin_{x\in\mathcal{X}}g(x|\lambda)\forall\lambda\in\Delta^{{m-1}}$, we need to randomly sample different trade-off preferences $\lambda\in\Delta^{{m-1}}$ (and hence different single-objective subproblems) at each iteration. In this case, it could be very hard to maintain extra sequences for each subproblem, since the trade-offs could be infinite and come in an online manner. Our proposed STCH method is simple, lightweight, and with good theoretical properties, and we hope that it can inspire more follow-up work on developing more efficient gradient-based methods (e.g., improved minmax optimization algorithms) for multi-objective optimization.
$\mu$ $\alpha$ $\epsilon_{HOMO}$ $\epsilon_{LUMO}$ $<R^{2}>$ ZPVE $U_{0}$ $U$ $H$ $G$ $c_{v}$ $\Delta_{p}\uparrow$
STL 0.062 0.192 58.82 51.95 0.529 4.52 63.69 60.83 68.33 60.31 0.069 0.00
TCH 0.266 0.401 107.1 151.6 5.922 13.2 166.7 167.5 168.1 162.0 0.206 -252.2
Smoothed-GDA 0.252 0.424 105.2 157.1 4.561 11.8 162.4 159.3 149.9 152.2 0.198 -218.8
STCH (Ours) 0.166 0.260 94.48 101.2 1.850 4.88 58.34 58.68 58.70 58.27 0.104 -58.14
We also implemented and tested smoothed-GDA on the QM9 dataset. According to the results above, smoothed-GDA achieves better overall performance than the original TCH scalarization, but is significantly outperformed by our proposed STCH scalarization method. We have also tried to tune the hyperparameters for smoothed-GDA, but do not obtain any significant improvement.
Appendix D Detailed Experiment Setting
D.1 Multi-Task Learning Problems
In this work, we implement our STCH scalarization method for MTL problems with the LibMTL library (Lin & Zhang, 2023), and follow the experiment settings of Lin et al. (2023a).
NYUv2 (Silberman et al., 2012) is an indoor scene understanding dataset with $3$ tasks on semantic segmentation, depth estimation, and surface normal prediction, with $795$ training and $654$ testing images. Following Lin et al. (2023a), we compare all MTL methods using the SegNet model with a shared encoder and 3 task-specific decoders. The model is trained for $200$ epochs with Adam (Kingma & Ba, 2015), of which the learning rate is initially set to $10^{-4}$ with $10^{-5}$ weight decay and will be halved to $5\times 10^{-5}$ after $100$ epochs. The batch size is set to $2$. The three optimization objectives are cross-entropy loss, $L_{1}$ loss and consine loss for semantic segmentation, depth estimation, and surface normal prediction.
Office-31 (Saenko et al., 2010) is an image classification dataset that contains $4,110$ images across $3$ domains (Amazon, DSLR, and Webcam), of which each task has $31$ classes. The data split from (Lin et al., 2022a) is utilized to split the data as $60\%$-$20\%$-$20\%$ for training, validation, and testing. All MTL methods are compared on the same model with a shared ResNet-18 pretrained encoder and linear task-specific head for each task. All methods are trained by Adam (Kingma & Ba, 2015), of which the learning rate is $10^{-4}$ with $10^{-5}$ weight decay. The batch size is $64$ and the training epoch is $100$. The optimization objectives are three cross-entropy losses for each classification task.
QM9 (Ramakrishnan et al., 2014) is a molecular property prediction dataset with 11 tasks on different properties. The data split in Navon et al. (2022) is used to divide the dataset into $110,000$ for training, $10,000$ for validation, and $10,000$ for testing. All MTL methods are compared on the same model with a shared graph neural network encoder and $11$ linear task-specific head as in Navon et al. (2022) and Lin et al. (2023a). All methods are trained by Adam (Kingma & Ba, 2015), of which the learning rate is $10^{-3}$ with the ReduceLROnPlateau scheduler. The batch size is $128$ and the number of training epochs is $300$. The optimization objectives are $11$ mean squared error (MSE) for each task.
D.2 Pareto Set Learning
Following other related works on Pareto set learning, we build a simple fully-connected multi-layered perceptron (MLP) as the Pareto set model:
$h_{\bm{\theta}}(\bm{\lambda}):$ Input ($\text{Input }(\bm{\lambda})\rightarrow\text{Linear}(m,256)\rightarrow\text{ReLU}\rightarrow\text{Linear}(256,256)$
$\rightarrow\text{ReLU}\rightarrow\text{Linear}(256,256)\rightarrow\text{ReLU}$ (46)
$\rightarrow\text{Linear}(256,n)\rightarrow\text{Output }\bm{x}(\bm{\lambda}),$
where the input is the preference $\bm{\lambda}\in\Delta^{m-1}$ with size $m$. This model is a two-layer MLP that has $256$ units in each hidden layer with ReLU activation, and the output is a solution $\bm{x}(\bm{\lambda})\in\mathbb{R}^{n}$. For Pareto set learning, the goal is to find the optimal model parameter $\bm{\theta}^{*}$ such that $\bm{x}^{*}(\bm{\lambda})=h_{\bm{\theta}^{*}}(\bm{\lambda})$ is the corresponding optimal solution for any given preference $\bm{\lambda}$. In all experiments, we train each set model with $2,000$ iterations, of which $10$ different preferences are uniformly sampled from $\Delta^{m-1}$ at each iteration. In other words, the total evaluation budget is $20,000$ for each method.
D.2.1 Synthetic Benchmark Problems
We first compare different scalarization methods for learning the Pareto set on $6$ synthetic benchmark problems shown on the next page. The problems F1-F3 have convex Pareto fronts, and the rest problems F4-F6 have concave Pareto fronts.
F1 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\sqrt{\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (47)
F2 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\sqrt{\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (48)
F3 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\sqrt{\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (49)
F4 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\left(\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}\right)^{2}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-(2\bm{x}_{1}-1)^{2})^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (50)
F5 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\left(\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}\right)^{2}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\bm{x}_{1}^{0.5(1.0+\frac{3(j-2)}{n-2})})^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (51)
F6 $f_{1}(\bm{x})=(1+\frac{s_{1}}{|J_{1}|})\bm{x}_{1},\quad f_{2}(\bm{x})=(1+\frac{s_{2}}{|J_{2}|})\left(1-\left(\frac{\bm{x}_{1}}{1+\frac{s_{2}}{|J_{2}|}}\right)^{2}\right)$ where $\text{where }s_{1}=\sum_{j\in J_{1}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2}\text{ and }s_{2}=\sum_{j\in J_{2}}(\bm{x}_{j}-\sin(4\pi\bm{x}_{1}+\frac{j\pi}{n}))^{2},$ $J_{1}=\{j|j\text{ is odd and }2\leq j\leq n\}\text{ and }J_{2}=\{j|j\text{ is even and }2\leq j\leq n\}$ (52)
D.2.2 Real-World Engineering Design Problems
In addition to synthetic benchmark problems, we also compare the STCH scalarization with different methods on the following five real-world engineering design problems summarized in Tanabe & Ishibuchi (2020):
Four Bar Truss Design
It is a multi-objective engineering design problem for a four-bar truss system proposed in Cheng & Li (1999):
$f_{1}(\bm{x})=L(2\bm{x}_{1}+\sqrt{2}\bm{x}_{2}+\sqrt{\bm{x}_{3}}+\bm{x}_{4})$ (53)
$f_{2}(\bm{x})=\frac{FL}{E}(\frac{2}{\bm{x}_{1}}+\frac{2\sqrt{2}}{\bm{x}_{2}}-\frac{2\sqrt{2}}{\bm{x}_{3}}+\frac{2}{\bm{x}_{4}})$
where the first objective $f_{1}(\bm{x})$ is to minimize the structural volume and the second objective $f_{2}(\bm{x})$ is to minimize the joint displacement. The four decision variables $\bm{x}_{1},\bm{x}_{4}\in[a,3a]$, $\bm{x}_{2},\bm{x}_{3}\in[\sqrt{2}a,3a]$ are the lengths of four bars with $a=F/\sigma$, and we have $F=10$, $E=2\times 10^{5}$, $L=200$, and $\sigma=10$.
Hatch Cover Design
It is a multi-objective engineering design problem for a hatch cover proposed in Amir & Hasegawa (1989):
$f_{1}(\bm{x})=\bm{x}_{1}+120\bm{x}_{2}$ (54)
$f_{2}(\bm{x})=\sum_{1=1}^{4}\max\{-g_{i}(\bm{x}),0\}\text{ where}$
$g_{1}(\bm{x})=1.0-\frac{\sigma_{b}}{\sigma_{b,\max}},$
$g_{2}(\bm{x})=1.0-\frac{\tau}{\tau_{\max}},$
$g_{3}(\bm{x})=1.0-\frac{\delta}{\delta_{\max}},$
$g_{4}(\bm{x})=1.0-\frac{\sigma_{b}}{\sigma_{k}}$
[…]
Disk Brake Design
It is a multi-objective engineering design problem for a disk brake proposed in Ray & Liew (2002):
$f_{1}(\bm{x})=4.9\times 10^{-5}(\bm{x}_{2}^{2}-\bm{x}_{1}^{2})(\bm{x}_{4}-1)$ (55)
$f_{2}(\bm{x})=9.82\times 10^{6}(\frac{\bm{x}_{2}^{2}-\bm{x}_{1}^{2}}{\bm{x}_{3}\bm{x}_{4}(\bm{x}_{2}^{3}-\bm{x}_{1}^{3})})$
$f_{3}(\bm{x})=\sum_{1=1}^{4}\max\{-g_{i}(\bm{x}),0\}\text{ where}$
$g_{1}(\bm{x})=(\bm{x}_{2}-\bm{x}_{1})-20,$
$g_{2}(\bm{x})=0.4-\frac{\bm{x}_{3}}{3.14(\bm{x}_{2}^{2}-\bm{x}_{1}^{2})},$
$g_{3}(\bm{x})=1-\frac{2.22\times 10^{-3}\bm{x}_{3}(\bm{x}_{2}^{3}-\bm{x}_{1}^{3})}{(\bm{x}_{2}^{2}-\bm{x}_{1}^{2})^{2}},$
$g_{4}(\bm{x})=\frac{2.66\times 10^{-2}\bm{x}_{3}\bm{x}_{4}(\bm{x}_{2}^{3}-\bm{x}_{1}^{3})}{(\bm{x}_{2}^{2}-\bm{x}_{1}^{2})}-900$
where the first objective $f_{1}(\bm{x})$ is the mass of the brake, the second objective $f_{2}(\bm{x})$ is the minimum stopping time of the disc brake, and the third objective is the sum of four constraint violations $g_{1}(\bm{x})$, $g_{2}(\bm{x})$, $g_{3}(\bm{x})$ and $g_{4}(\bm{x})$. For the decision variables, $\bm{x}_{1}\in[55,80]$ is the inner radius of the discs, $\bm{x}_{2}\in[75,110]$ is the outer radius of the discs, $\bm{x}_{3}\in[1000,3000]$ is the engaging force, and $\bm{x}_{4}\in[11,20]$ is the number of friction surfaces.
Gear Train Design
It is a multi-objective engineering design problem for a gear train proposed in Deb & Srinivasan (2006):
$f_{1}(\bm{x})=|6.931-\frac{\bm{x}_{3}\bm{x}_{4}}{\bm{x}_{1}\bm{x}_{2}}|$ (56)
$f_{2}(\bm{x})=\max\{\bm{x}_{1},\bm{x}_{2},\bm{x}_{3},\bm{x}_{4}\}$
$f_{3}(\bm{x})=\max\{-g_{1}(\bm{x}),0\}\text{ where}$
$g_{1}(\bm{x})=0.5-\frac{f_{1}(\bm{x})}{6.931}.$
The first objective $f_{1}(\bm{x})$ is the difference between the realized gear ration and a given specific gear ration, the second objective $f_{2}(\bm{x})$ is the maximum size of the four gears, and the third objective $f_{3}(\bm{x})$ is the constraint violation of $g_{1}(\bm{x})$. The four integer decision variables $\bm{x}_{i}\in\{12,\ldots,60\}$ are the number of teeth in each of the four gears.
Rocket Injector Design
It is a multi-objective engineering design problem for a rocket injector proposed in Vaidyanathan et al. (2003):
$f_{1}(\bm{x})=$ $0.692+0.477\bm{x}_{1}-0.687\bm{x}_{2}-0.080\bm{x}_{3}-0.0650\bm{x}_{4}$ (57)
$-0.167\bm{x}_{1}\bm{x}_{1}-0.0129\bm{x}_{2}\bm{x}_{1}+0.0796\bm{x}_{2}\bm{x}_{2}-0.00634\bm{x}_{3}\bm{x}_{1}-0.0257\bm{x}_{3}\bm{x}_{2}$
$+0.0877\bm{x}_{3}\bm{x}_{3}-0.0521\bm{x}_{4}\bm{x}_{1}+0.00156\bm{x}_{4}\bm{x}_{2}+0.00198\bm{x}_{4}\bm{x}_{3}+0.0184\bm{x}_{4}\bm{x}_{4}$
$f_{2}(\bm{x})=$ $0.153+0.322\bm{x}_{1}-0.396\bm{x}_{2}-0.424\bm{x}_{3}-0.0226\bm{x}_{4}$
$-0.175\bm{x}_{1}\bm{x}_{1}-0.0185\bm{x}_{2}\bm{x}_{1}+0.0701\bm{x}_{2}\bm{x}_{2}-0.251\bm{x}_{3}\bm{x}_{1}-0.179\bm{x}_{3}\bm{x}_{2}$
$+0.0150\bm{x}_{3}\bm{x}_{3}-0.0134\bm{x}_{4}\bm{x}_{1}+0.0296\bm{x}_{4}\bm{x}_{2}+0.0752\bm{x}_{4}\bm{x}_{3}+0.0192\bm{x}_{4}\bm{x}_{4}$
$f_{3}(\bm{x})=$ $0.370+0.205\bm{x}_{1}-0.0307\bm{x}_{2}-0.108\bm{x}_{3}-1.019\bm{x}_{4}$
$-0.135\bm{x}_{1}\bm{x}_{1}-0.0141\bm{x}_{2}\bm{x}_{1}+0.0998\bm{x}_{2}\bm{x}_{2}-0.208\bm{x}_{3}\bm{x}_{1}-0.0301\bm{x}_{3}\bm{x}_{2}$
$+0.226\bm{x}_{3}\bm{x}_{3}-0.353\bm{x}_{4}\bm{x}_{1}+0.0497\bm{x}_{4}\bm{x}_{3}+0.423\bm{x}_{4}\bm{x}_{4}+0.202\bm{x}_{2}\bm{x}_{1}\bm{x}_{1}$
$-0.281\bm{x}_{3}\bm{x}_{1}\bm{x}_{1}-0.342\bm{x}_{2}\bm{x}_{2}\bm{x}_{1}-0.245\bm{x}_{2}\bm{x}_{2}\bm{x}_{3}+0.281\bm{x}_{3}\bm{x}_{3}\bm{x}_{2}$
$-0.184\bm{x}_{4}\bm{x}_{4}\bm{x}_{1}-0.281\bm{x}_{2}\bm{x}_{1}\bm{x}_{3}$
where the first objective $f_{1}(\bm{x})$ is the maximum temperature of the injector face, the second objective $f_{2}(\bm{x})$ is the distance from the inlet, and the third objective $f_{3}(\bm{x})$ is the maximum temperature at the post tip. The decision variables $\bm{x}=[\bm{x}_{1},\bm{x}_{2},\bm{x}_{3},\bm{x}_{4}]\in[0,1]^{4}$ are the hydrogen flow angle ($\alpha$), the hydrogen area ($\Delta$HA), the oxygen area ($\Delta$OA), and the oxidizer post tip thickness (OPTT), respectively.
D.2.3 Hypervolume Definition
Following the related works on multi-objective optimization, we use the hypervolume metrics (Zitzler et al., 2007) to compare the qualities of approximate Pareto sets obtained by different methods. For a solution set $P$, its hypervolume $\text{HV}(P)$ in the objective space can be defined as the volume of the following dominated set:
$S=\{\bm{r}\in\mathbb{R}^{m}\mid\exists\bm{y}\in P\mbox{ such that }\bm{y}\prec\bm{r}\prec\bm{r}^{*}\},$ (58)
where $\bm{r}^{*}$ is the reference point that is dominated by all solutions in $P$, every point $\bm{r}\in S$ will be dominated by at least one $\bm{y}\in P$, and we have $\text{HV}(P)=\textbf{Vol}(S)$. According to this definition, if a solution set $A$ dominates another solution set $B$ (e.g., every $\bm{b}\in B$ is dominated by at least one $\bm{a}\in A$), we will always have $\text{HV}(A)>\text{HV}(B)$. The ground-truth Pareto set $P^{*}$ will always have the largest hypervolume $\text{HV}(P^{*})$. In the experiments, we report the hypervolume difference to the optimal Pareto set for each algorithm:
$\Delta\text{HV}(P)=\text{HV}(P^{*})-\text{HV}(P)$ (59)
where a solution set $P$ with better quality should have smaller $\Delta\text{HV}(P)$ and $\Delta\text{HV}(P^{*})=0$. We report the average $\Delta\text{HV}(P)$ over $30$ independent runs for each method.
Appendix E Additional Experimental Studies
E.1 Office-31
Table 6: Results on the Office-31 dataset.
[…]
Single-Task Baseline
[…]
Adaptive Gradient Method
MGDA 85.47 95.90 97.03 92.80$\pm$0.14 -0.27$\pm$0.15
GradNorm 83.58 97.26 96.85 92.56$\pm$0.87 -0.59$\pm$0.94
PCGrad 83.59 96.99 96.85 92.48$\pm$0.53 -0.68$\pm$0.57
GradDrop 84.33 96.99 96.30 92.54$\pm$0.42 -0.59$\pm$0.46
GradVac 83.76 97.27 96.67 92.57$\pm$0.73 -0.58$\pm$0.78
IMTL-G 83.41 96.72 96.48 92.20$\pm$0.89 -0.97$\pm$0.95
CAGrad 83.65 95.63 96.85 92.04$\pm$0.79 -1.14$\pm$0.85
MTAdam 85.52 95.62 96.29 92.48$\pm$0.87 -0.60$\pm$0.93
Nash-MTL 85.01 97.54 97.41 93.32$\pm$0.82 +0.24$\pm$0.89
MetaBalance 84.21 95.90 97.40 92.50$\pm$0.28 -0.63$\pm$0.30
[…]
Aligned-MTL 83.36 96.45 97.04 92.28$\pm$0.46 -0.90$\pm$0.48
IMTL 83.70 96.44 96.29 92.14$\pm$0.85 -1.02$\pm$0.92
[…]
Adaptive Loss Method
UW 83.82 97.27 96.67 92.58$\pm$0.84 -0.56$\pm$0.90
DWA 83.87 96.99 96.48 92.45$\pm$0.56 -0.70$\pm$0.62
IMTL-L 84.04 96.99 96.48 92.50$\pm$0.52 -0.63$\pm$0.58
[…]
Scalarization Method
EW 83.53 97.27 96.85 92.55$\pm$0.62 -0.61$\pm$0.67
GLS 82.84 95.62 96.29 91.59$\pm$0.58 -1.63$\pm$0.61
RLW 83.82 96.99 96.85 92.55$\pm$0.89 -0.59$\pm$0.95
TCH 84.45 96.72 96.11 92.43$\pm$0.46 -0.71$\pm$0.56
[…]
Full results on the Office-31 dataset are shown in Table 6. We run the experiments with MoCO, TCH, and STCH methods by ourselves, and the rest results are from Lin et al. (2023a) with the same setting. Our proposed STCH scalarization method performs the best on the Amazon task, and achieves the second best performance on DSLR and Webcam, which leads to the best average performance and the best $\Delta_{p}$. In fact, it is the only method that can dominate the STL baseline on all tasks.
E.2 QM9
Table 7: Results on the QM9 dataset.
$\mu$ $\alpha$ $\epsilon_{HOMO}$ $\epsilon_{LUMO}$ $\langle R^{2}\rangle$ ZPVE $U_{0}$ $U$ $H$ $G$ $c_{v}$ $\Delta_{p}\uparrow$
Single-Task Baseline
STL 0.062 0.192 58.82 51.95 0.529 4.52 63.69 60.83 68.33 60.31 0.069 0.00
Adaptive Gradient Method
MGDA 0.181 0.325 118.6 92.45 2.411 5.55 103.7 104.2 104.4 103.7 0.110 -103.0$\pm$8.62
[…]
PCGrad 0.104 0.293 75.29 88.99 3.695 8.67 115.6 116.0 116.2 113.8 0.109 -117.8$\pm$3.97
GradDrop 0.114 0.349 75.94 94.62 5.315 15.8 155.2 156.1 156.6 151.9 0.136 -191.4$\pm$9.62
GradVac 0.100 0.299 68.94 84.14 4.833 12.5 127.3 127.8 128.1 124.7 0.117 -150.7$\pm$7.41
IMTL-G 0.670 0.978 220.7 249.7 19.48 55.6 1109 1117 1123 1043 0.392 -1250$\pm$90.9
CAGrad 0.107 0.296 75.43 88.59 2.944 6.12 93.09 93.68 93.85 92.32 0.106 -87.25$\pm$1.51
MTAdam 0.593 1.352 232.3 419.0 24.31 69.7 1060 1067 1070 1007 0.627 -1403$\pm$203
[…]
MoCo 0.489 1.096 189.5 247.3 34.33 64.5 754.6 760.1 761.6 720.3 0.522 -1314$\pm$65.2
Aligned-MTL 0.123 0.295 98.07 94.56 2.397 5.90 86.42 87.42 87.19 86.75 0.106 -80.58$\pm$4.18
IMTL 0.138 0.344 106.1 102.9 2.595 7.84 102.5 103.0 103.2 100.8 0.110 -104.3$\pm$11.7
DB-MTL 0.112 0.264 89.26 86.59 2.429 5.41 60.33 60.78 60.80 60.59 0.098 -58.10$\pm$3.89
Adaptive Loss Method
UW 0.336 0.382 155.1 144.3 0.965 4.58 61.41 61.79 61.83 61.40 0.116 -92.35$\pm$13.9
DWA 0.103 0.311 71.55 87.21 4.954 13.1 134.9 135.8 136.3 132.0 0.121 -160.9$\pm$16.7
[…]
IGBv2 0.235 0.377 132.3 139.9 2.214 5.90 64.55 65.06 65.12 64.28 0.121 -99.86$\pm$10.4
Scalarization Method
EW 0.096 0.286 67.46 82.80 4.655 12.4 128.3 128.8 129.2 125.6 0.116 -146.3$\pm$ 7.86
GLS 0.332 0.340 143.1 131.5 1.023 4.45 53.35 53.79 53.78 53.34 0.111 -81.16$\pm$15.5
RLW 0.112 0.331 74.59 90.48 6.015 15.6 156.0 156.8 157.3 151.6 0.133 -200.9$\pm$13.4
TCH 0.266 0.401 107.1 151.6 5.922 13.2 166.7 167.5 168.1 162.0 0.206 -252.2$\pm$16.6
[…]
Full results on the QM9 dataset are shown in Table 7. We run the experiments with TCH and STCH by ourselves, and the other results are from Lin et al. (2023a). Our STCH scalarization has a promising second best overall performance, while the gap to the best DB-MTL (Lin et al., 2023a) is very tight. It should be noticed that DB-MTL is an adaptive gradient method that requires a much longer runtime as reported in Table 4. Therefore, our proposed STCH scalarization method can serve as a very promising alternative for solving MTL problems.
E.3 More Comparison to Linear Scalarization
Table 8: Results on the CelebA dataset.
Average Task Accuracy
Unit. Scal. 9.090e-01$\pm$7.568e-04
IMTL 9.093e-01$\pm$7.631e-04
MGDA 9.022e-01$\pm$9.687e-04
GradDrop 9.098e-01$\pm$3.383e-04
PCGrad 9.093e-01$\pm$1.108e-03
RLW Diri. 9.099e-01$\pm$7.845e-04
RLW Norm. 9.095e-01$\pm$1.012e-03
STCH (Ours) 9.098e-01$\pm$4.692e-04
We also test our proposed STCH scalarization on the CelebA dataset with the same setting as in Kurin et al. (2022). We independently run STCH $3$ times and report the mean of average task accuracy and standard deviation, while all other results are directly from Kurin et al. (2022). STCH has a simple equal weight for all tasks which is not tuned. According to the results in Table 8, STCH achieves similar performance with most other methods, and significantly outperforms MGDA. We find that STCH achieves almost zero training loss for all tasks, which is consistent with most other methods as reported in Kurin et al. (2022). In this case, the testing accuracy mainly depends on the generalization performance rather than the optimization performance (on training loss). Since the STCH scalarization is developed from the viewpoint of multi-objective optimization, we currently do not have a straightforward way to analyze its generalization performance. Recently, Chen et al. (2023) have investigated the generalization gap for the stochastic adaptive gradient method. We leave the generalization analysis for STCH in the stochastic setting to future work.
On the other hand, our proposed STCH scalarization is lightweight, with a computational overhead similar to that of linear scalarization. It is also interesting to further investigate its performance with tuned weights and/or regularization for multi-task learning as in Kurin et al. (2022) and Xin et al. (2022).
E.4 Multi-Objective Bayesian Optimization
Table 9: Results (hypervolume difference $\Delta\text{HV}\downarrow$) for multi-objective Bayesian optimization on $6$ synthetic benchmark problems and $5$ real-world engineering design problems.
F1 F2 F3 F4 F5 F6 BarTruss PressureVessel DiskBrake GearTrain RocketInjector
TCH 0.0183 0.0185 0.0413 0.0891 0.0602 0.0917 0.0429 0.0441 0.0262 0.0262 0.0449
STCH 0.0147 0.0110 0.0304 0.0476 0.0455 0.0575 0.0233 0.0194 0.0214 0.0198 0.0332
In this subsection, we test our proposed STCH on multi-objective Bayesian optimization. We follow the setting proposed in Lin et al. (2022c), where a Pareto set learning approach with TCH scalarization (PSL-TCH) has been proposed for multiobjective Bayesian optimization. We simply replace the TCH scalarization in PSL-TCH with the proposed STCH scalarization to obtain a new method PSL-STCH. We then compare the new PSL-STCH with PSL-TCH on different synthetic and multiobjective engendering design problems with a small evaluation budget (i.e. 100 evaluations) as in Lin et al. (2022c). According to Table 9 with the hypervolume difference ($\Delta HV$) metrics, PSL-STCH can outperform PSL-TCH on all problems. This result confirms that the proposed STCH scalarization can serve as an off-the-shelf approach for different multiobjective optimization applications, including Bayesian optimization.
E.5 Efficient Pareto Set Learning
[…]
F1 F2 F3 F4 F5 F6 BarTruss HatchCover DiskBrake GearTrain RocketInjector
LS 0.8602 0.8629 0.7826 0.3172 0.3712 0.2892 0.8805 1.1633 0.9738 1.0169 0.7051
COSMOS 0.8608 0.8614 0.8638 0.5283 0.5300 0.5242 0.8803 1.1425 0.9710 1.0174 0.8091
EPO 0.8653 0.8689 0.8564 0.5324 0.5349 0.5236 0.8772 1.1592 0.9805 1.0175 0.7889
TCH 0.8676 0.8686 0.8582 0.5345 0.5364 0.5287 0.8795 1.1611 0.9765 1.0170 0.8198
STCH 0.8707 0.8709 0.8670 0.5365 0.5372 0.5316 0.8829 1.1632 0.9864 1.0177 0.8363
Hypervolume Metrics
For the efficient Pareto set learning problem, we report the hypervolume for all problems in Table 10. For each problem, we first normalize the objective value of all solutions for all methods into $[0,1]$ with the same nadir point and ideal point, then calculate the hypervolume with reference point $[1.1,1.1]^{m}$ where $m$ is the number of objectives. Our proposed STCH achieves the best overall performance on these problems.
Table 11: The run time for learning the Pareto set of $5$ real-world enginnering design problems with different methods.
Method LS COSMOS EPO TCH STCH
Bar Truss 6s (1x) 6s (1x) 268s (45x) 6s (1x) 6s (1x)
Hatch Cover 6s (1x) 6s (1x) 277s (46x) 6s (1x) 6s (1x)
Disk Brake 9s (1x) 9s (1x) 329s (37x) 9s (1x) 9s (1x)
Gear Train 9s (1x) 9s (1x) 337s (38x) 9s (1x) 9s (1x)
Rocket Injector 9s (1x) 9s (1x) 324s (36x) 9s (1x) 9s (1x)
Run time
In addition, we also report the runtimes of different methods to learn the Pareto set for five multi-objective engineering design problems in Table 11. According to the results, our proposed STCH scalarization method share a similar runtime with other scalarization methods (e.g., LS, COSMOS, and TCH), while the gradient-based EPO method takes significantly longer runtime ($36$x to $44$x) to learn the Pareto set.
[…]
Figure 6: The log hypervolume difference ($\log(\Delta HV)$) of different methods during the optimization process for $11$ different multi-objective optimization problems.